### v1

# MAVERIX: Multimodal Audio-Visual Evaluation Reasoning IndeX

Abstract
Frontier models have either been language-only or have primarily focused on vision and language modalities. Although recent advancements in models with vision and audio understanding capabilities have shown substantial progress, the field lacks a standardized evaluation framework for thoroughly assessing their cross-modality perception performance. We introduce MAVERIX (Multimodal Audio-Visual Evaluation Reasoning IndeX), a novel benchmark with 700 videos and 2,556 questions explicitly designed to evaluate multimodal models through tasks that necessitate close integration of video and audio information. MAVERIX uniquely provides models with audiovisual tasks, closely mimicking the multimodal perceptual experiences available to humans during inference and decision-making processes. To our knowledge, MAVERIX is the first benchmark aimed explicitly at assessing comprehensive audiovisual integration. Experiments with state-of-the-art models, including Gemini 1.5 Pro and o1, show performance approaching human levels (around 70% accuracy), while human experts reach near-ceiling performance (95.1%). With standardized evaluation protocols, a rigorously annotated pipeline, and a public toolkit (maverix-benchmark.github.io) [^1] Project page will be available soon., MAVERIX establishes a challenging testbed for advancing audiovisual multimodal intelligence.
Figure 1: MAVERIX contains over 2,500 audio-visual tasks spanning diverse domains with tight modality interdependence (left) and remains challenging, as state-of-the-art multimodal frontier models are consistently outperformed by average human participants (right).
Figure 2: Example questions and answers in the MAVERIX benchmark.
1 Introduction
Human cognition seamlessly integrates visual and auditory information to reason, infer, and interact within dynamic environments. Replicating this ability in Multimodal Large Language Model (MLLM) systems remains a central challenge for AI, as autonomous agents must process complex audiovisual input to engage meaningfully with the world [32, 4, 1].
Recent progress in multimodal foundation models has brought us closer to this goal, but current benchmarks fall short in assessing their abilities to reason with multimodal inputs. Most benchmarks focus on static images [11, 3], simple recognition, or tasks that can be solved through unimodal cues like transcripts [44, 13]. These benchmarks fail to probe the deeper, joint reasoning across modalities needed for real-world scenarios, such as interpreting social interactions or anticipating off-screen events [10, 36].
A core obstacle in designing effective multimodal benchmarks is ensuring that tasks genuinely require multimodal understanding rather than allowing models to exploit unimodal shortcuts. Many existing video-language benchmarks reduce to visible-object recognition or dialogue parsing, sidestepping the need to synthesize audio-visual dependencies [40, 25, 29].
To address this, we introduce MAVERIX, a benchmark designed to evaluate multimodal video-audio understanding through tasks that have tight modality interdependence. MAVERIX features four challenging tasks: situational awareness, contextual summarization, social sentiment analysis, and dynamic reasoning, covering 700 videos and over 2,500 carefully designed questions. These are constructed through a hybrid human-AI pipeline to ensure that solving them requires intertwined audiovisual synthesis, revealing the underlying multimodal reasoning capabilities of models applied to the benchmark tasks.
Evaluations of state-of-the-art models, including Gemini 1.5 Pro [49], GPT-4o [39], and o1 [38], reveal significant gaps. Gemini 1.5 Pro, even with direct audio-video inputs, achieves 71.9% accuracy on multiple choice questions, marginally lower than human performance (80%). Open-ended responses further expose weaknesses in temporal reasoning and contextual understanding, with models averaging only 1.9/5 vs. human 2.79/5 in GPT-4o-judged scoring. Further, models that are not capable of processing raw audio and rely solely on transcripts perform even worse, highlighting the inadequacy of text-only proxies for rich audiovisual comprehension [37, 17].
By providing a unified evaluation framework, high-quality human-validated questions, and an open-source toolkit, MAVERIX aims to advance research toward robust multimodal reasoning at the human level.
2 Benchmark Design and Constructions
Designing benchmarks for audiovisual understanding requires tasks that demand genuine cross-modal reasoning, not just unimodal pattern matching. In real-world scenarios, one modality often dominates, such as visuals in a silent event or audio in a podcast, but effective understanding depends on integrating cues across modalities. Existing benchmarks typically focus on narrow tasks or merge disjointed datasets, limiting their ability to evaluate general-purpose multimodal LLMs (MLLMs). Moreover, many video-language benchmarks rely on multiple-choice questions (MCQs) for their simplicity, but small answer spaces of the common four-option MCQs inflate scores through random guessing and fail to assess free-form reasoning — a core aspect of human-like comprehension.
Figure 3: Annotation workflow outlining the stages of data selection, review, and refinement.
To address these limitations, MAVERIX introduces a unified benchmark of diverse real-world tasks that systematically test the interdependence of modality while supporting a dual evaluation framework of eight-option MCQs and open-ended questions. The expanded eight-option format reduces guess bias (12.5% baseline), providing more reliable accuracy metrics, while open-ended tasks require models to generate nuanced, context-aware responses. For open-ended evaluation, we employ GPT-4o as an automated judge, scoring responses against human-curated references using a five-dimensional rubric: correctness of information, attention to detail, contextual understanding, temporal coherence, and consistency under paraphrasing. MAVERIX’s taxonomy and revealed performance gaps are shown in Fig. 1.
This hybrid design balances scalability and depth: MCQs provide standardized large-scale evaluation, while open-ended questions probe the models’ ability to synthesize and express complex audiovisual reasoning. Together, they create a robust assessment ecosystem tailored to intertwined multimodal understanding. We detail our pipeline used to generate MAVERIX’s 2,500+ high-quality questions and answers.
2.1 Dataset Generation Pipeline
Dataset Collection To ensure broad coverage of multimodal reasoning, we designed four domain hierarchies spanning agentic tasks, topics, video categories, and multimodal abilities, as shown in Fig. 2. Under agentic tasks, we defined seven key areas—social interactions, sentiment expression, information querying, sports analysis, gaming commentary, egocentric navigation, and multilingual contexts—each reflecting real-world scenarios that require complementary audiovisual cues. For example, gaming videos involve coordinating visual gameplay with voice chat strategies. In total, we annotated 155 fine-grained activity tags (e.g., “referee decision analysis” under sports) to capture diverse reasoning challenges (see Fig. 5).
Figure 4: Overview of the dataset distribution. The corresponding number of samples in each category is shown.
We sourced videos from YouTube-8M [2], MSR-VTT [53], UR-FUNNY-V2 [20], Ego4D [19], and AudioSet [18], balancing short (<1absent1<1< 1 min), medium (2–10 min), and long (10–65 min) formats. Each video includes subtitles, categorized audio (e.g., speech, environmental, artificial), and visual keyframes sampled at 1-second intervals. After deduplication and quality checks, MAVERIX comprises 700 videos with balanced coverage across domains and durations.
Multimodal Question-Answer Annotation To evaluate models’ ability to synthesize audiovisual information, we created 852 eight-option multiple-choice questions through a three-stage annotation protocol. A team of 11 researchers proficient in English (or another second language) with expertise in vision-language learning ensured high-quality, semantically precise questions.
The process involved (1) selecting videos and identifying clips requiring understanding multiple modalities, (2) drafting questions designed to be answerable only through joint audiovisual analysis, and (3) applying dual-validation checks to eliminate unimodal shortcuts by verifying that neither audio-only nor vision-only playback sufficed. The correct option is manually annotated, then paraphrased and modified into semantically different options with structurally similar alternative options. Surviving questions underwent peer review, with ambiguous cases resolved by consensus.
Figure 5: Dataset statistics. Left: distribution of video sources. Center: question and answer length across difficulty levels. Right: distribution of video durations for full-length and truncated clips.
For example, a valid question might ask, “Why did the mechanic abruptly stop speaking?” requiring both visual cues (e.g., discovering a leak) and audio cues (e.g., sudden silence). This protocol ensures MAVERIX’s QA pairs demand genuine modality interdependence, preventing reliance on any single modality.
Quality Assurance To ensure the reliability of MAVERIX’s videos and annotations, we apply a two-stage quality control protocol combining expert review and model-based validation. Each QA pair undergoes four checks by a separate annotator, as illustrated in Fig. 3: (1) linguistic validity for clear and grammatical phrasing, (2) answerability of whether the question is resolvable via the video’s audiovisual content), (3) option integrity to ensure one correct answer with plausible distractors like semantically tangent or structurally identical options, and (4) modality interdependence, using cross-modal invalidation tests from Section 3.1 (e.g., disabling audio or video to detect shortcuts). For open-ended questions, reviewers also confirm that rephrased variants preserve meaning without overlapping with the ground-truth wording.
To guard against textual or visual biases, we test each question with GPT-4o for text-only and video keyframes + subtitles inputs. Questions answered correctly without full audiovisual reasoning are revised. For example, if GPT-4o answers “Why did the mechanic sigh?” from subtitles alone, we refine the question. Reviewers document model failures and work with annotators to iteratively resolve ambiguous cases. This combined human-model process ensures MAVERIX’s questions demand proper multimodal understanding.
2.2 MAVERIX Statistics
MAVERIX comprises a diverse set of videos spanning 155 real-world scenarios across seven agentic task domains, such as social interactions, sentiment analysis, and egocentric understanding. These are complemented by hierarchies over topics (e.g., travel, technology), video categories (e.g., documentaries, vlogs), and multimodal abilities (e.g., temporal reasoning, emotion recognition), supporting broad and balanced evaluation. Each question is also tagged with a difficulty level reflecting its multimodal complexity as judged by human annotators. A detailed breakdown of these splits is provided in Fig. 4.
The dataset contains 105.8 hours of video footage, with durations ranging from 10 seconds to 63 minutes. Videos are distributed across three duration categories: 16.8% short clips (<1 minute) for rapid context-switching tasks, 75.7% medium-length videos (1–10 minutes) for sustained narrative reasoning, and 7.4% long-form content (10–65 minutes) for testing temporal coherence. The average video length is 5.7 minutes, reflecting real-world use cases like troubleshooting tutorials or sports highlights.
Each video is paired with 3–4 questions on average (2,556 total), including eight-option multiple-choice questions (MCQs) and open-ended prompts. Questions are evenly rebalanced across tasks to ensure a balanced evaluation. To mitigate positional bias, answer labels are uniformly redistributed across options (e.g., correct answers appear equally in positions A–H). Fig. 4 details the dataset’s domain composition, and Fig. 5 shows the duration distribution and question-type breakdown.
3 Experiments
3.1 Evaluation Protocol
MAVERIX adopts a dual evaluation framework to assess multimodal LLMs (MLLMs) through eight-option multiple-choice questions (MCQs) and open-ended generation tasks.
Evaluation is conducted under two settings: localized, where models access only the timestamped video segment relevant to each question, and global, where the full-length video is provided. The localized setting mitigates information leakage by limiting the context to the specific temporal window required for reasoning.
For MCQs, we report both task-specific and overall accuracy, with answer choices uniformly distributed across positions (A–H) to reduce positional bias. Open-ended responses are evaluated using a GPT-4o-based scoring pipeline, adapted from Video-ChatGPT [35], which assesses outputs on five dimensions using a 0–5 scale. Scoring is calibrated on 200 human-annotated examples to ensure consistency, with penalties applied for failures in audiovisual reasoning, such as overlooking modality contradictions.
Results are aggregated across tasks (Tab. 1), with separate analyses for easy, medium, and hard videos to diagnose reasoning limitations. This protocol ensures reproducibility while addressing modality interdependence, which is the core challenge MAVERIX is designed to evaluate.
3.2 Baselines
We evaluate MAVERIX on a diverse suite of 12 MLLMs, encompassing both commercial and open-source architectures, to assess their ability to reason over intertwined audiovisual modalities. Commercial models include Gemini 1.5 Pro [49], GPT-4o [39], o1 [38], Grok2-Vision [52], Claude Sonnet 3.5 [5], Nova-Lite [23], and Nova-Pro [23], while open source representatives feature InternVL2 [14], Qwen2.5-VL [6], LLaVA-OneVision [27], DeepSeek-VL2-Tiny [51], and DeepSeek-VL2-Small [51]. Gemini 1.5 Pro is uniquely equipped with native audiovisual processing, enabling direct ingestion of raw video-audio streams. For these models, we maximize temporal resolution by sampling frames at their maximum supported rates. However, most architectures, including GPT-4o and LLaVA-OneVision, require transcribed subtitles as text proxies for audio. To standardize inputs, we preprocess all videos using Whisper-base-v3 to extract time-synced subtitles, with the timestamps provided to the evaluated models.
All models receive inputs in the unified format [video frames, subtitles, question], with frames uniformly sampled at their maximum supported context window. For audio-incapable models, subtitles replace raw audio tracks, while Gemini 1.5 Pro processes synchronized audio-video pairs. We employ a standardized prompt template across models, ensuring fairness by eliminating instructional biases. This setup isolates modality interdependence as the critical challenge: models must synthesize visual cues like the startled expression of a character with an auditory context like an off-screen crash to match human-like reasoning.
3.3 Results
Models Params Easy (%) Medium (%) Hard (%) Overall (%)
w/o sub w/ sub w/o sub w/ sub w/o sub w/ sub w/o sub w/ sub
Random Guess - 12.4 12.7 12.3 12.5
Human (Video) - 60.2 48.8 59.4 55.3
Human (Audio) - 34.9 27.4 40.8 31.1
Human (Video+Audio) - 83.6 76.2 84.4 80.3
Open-source MLLMs (Truncated Videos)
InternVL2 [14] 2B 31.1 35.2 (+4.1) 27.7 35.1 (+7.4) 17.8 24.3 (+6.5) 27.5 33.4 (+5.9)
Qwen2.5-VL [6] 7B 30.8 34.9 (+4.1) 25.9 31.9 (+6.0) 19.3 20.7 (+1.4) 26.8 31.3 (+4.5)
LLaVA-OneVision [27] 7B 26.6 28.9 (+2.3) 26.5 28.2 (+1.7) 29.9 29.9 (+0.0) 27.1 28.8 (+1.7)
DeepSeek-VL2-Tiny [51] 3.37B 16.8 15.1 (-1.7) 14.5 20.6 (+6.1) 20.5 17.3 (-3.2) 16.4 17.9 (+1.5)
DeepSeek-VL2-Small [51] 16.1B 39.9 46.4 (+6.5) 38.8 46.4 (+7.6) 25.9 29.6 (+3.7) 37.2 43.8 (+6.6)
Closed-source MLLMs (Truncated Videos)
GPT-4o [39] - 58.9 69.8 (+10.9) 60.7 67.0 (+6.3) 41.5 49.6 (+8.1) 56.9 65.4 (+8.5)
o1 [38] - 74.0 79.3 (+5.3) 70.7 76.0 (+5.3) 55.6 67.4 (+11.8) 69.6 75.9 (+6.3)
Claude Sonnet 3.5 [5] - 27.4 68.9 (+41.5) 39.6 64.6 (+25.0) 27.4 44.4 (+17.0) 40.3 63.1 (+22.8)
Grok2-Vision [52] - 58.3 65.1 (+6.8) 55.1 61.2 (+6.1) 45.9 47.4 (+1.5) 54.9 60.6 (+5.7)
Nova-Lite [23] - 51.8 59.8 (+8.0) 49.3 55.9 (+6.6) 37.8 43.0 (+5.2) 48.5 55.4 (+6.9)
Nova-Pro [23] - 53.0 51.5 (-1.5) 51.7 47.5 (-4.2) 35.6 34.1 (-1.5) 49.6 46.9 (-2.7)
Gemini 1.5 Pro [49] - 64.5 69.5 (+5.0) 65.4 69.4 (+4.0) 55.6 49.6 (-6.0) 63.5 66.3 (+2.8)
Open-source MLLMs (Full-Length Videos)
InternVL2 [14] 2B 23.0 26.3 (+3.3) 18.9 24.4 (+5.5) 14.2 11.0 (-3.2) 19.7 23.0 (+3.3)
Qwen2.5-VL [6] 7B 47.4 55.9 (+8.5) 45.3 51.7 (+6.4) 37.9 47.2 (+9.3) 44.8 52.6 (+7.8)
LLaVA-OneVision [27] 7B 34.5 40.1 (+5.6) 34.3 39.5 (+5.2) 30.7 35.4 (+4.7) 33.8 39.1 (+5.3)
DeepSeek-VL2-Tiny [51] 3.37B 17.1 26.0 (+8.9) 15.4 21.5 (+6.1) 21.3 19.7 (-1.6) 17.0 23.0 (+6.0)
DeepSeek-VL2-Small [51] 16.1B 28.3 44.1 (+15.8) 29.9 41.6 (+11.7) 23.6 26.0 (+2.4) 28.3 40.0 (+11.7)
Closed-source MLLMs (Full-Length Videos)
GPT-4o [39] - 57.6 64.8 (+7.2) 58.4 62.2 (+3.8) 43.3 46.5 (+3.2) 55.6 60.6 (+5.0)
o1 [38] - 72.4 73.0 (+0.6) 68.3 72.4 (+4.1) 52.0 65.4 (+13.4) 67.2 71.5 (+4.3)
Claude Sonnet 3.5 [5] - 7.9 17.4 (+9.5) 8.1 14.8 (+6.7) 7.9 17.3 (+9.4) 8.0 16.3 (+8.3)
Grok2-Vision [52] - 55.6 59.2 (+3.6) 55.8 55.2 (-0.6) 45.7 53.5 (+7.8) 54.1 56.5 (+2.4)
Nova-Lite [23] - 47.4 59.5 (+12.1) 45.3 55.9 (+10.6) 37.8 43.0 (+5.2) 44.9 55.3 (+10.4)
Nova-Pro [23] - 45.2 58.2 (+13.0) 48.4 58.1 (+9.7) 28.9 41.7 (+12.8) 44.1 55.5 (+11.4)
Gemini 1.5 Pro [49] - 65.1 69.4 (+4.3) 64.0 68.0 (+4.0) 54.3 55.1 (+0.8) 62.8 66.5 (+3.7)
Table 1: Performance comparison of different MLLMs across various settings with and without subtitles.
Video only Audio only Text only Video + Sub Video + Audio
Truncated Full Truncated Full Truncated Full Truncated Full Truncated Full
Level Easy 64.5 65.1 56.8 57.6 45.6 51.6 69.5 69.4 76.0 74.0
Medium 65.4 64.0 57.8 52.3 47.0 46.8 69.4 68.0 71.5 70.3
Hard 55.6 54.3 46.7 49.6 34.8 41.7 49.6 55.1 63.0 62.2
Agentic Tasks Social 55.6 58.3 47.5 49.0 41.4 54.2 62.6 59.4 61.6 67.7
Sentiment 69.1 70.4 53.1 51.9 34.6 42.0 66.7 64.2 76.5 74.1
Egocentric Agent 64.4 61.7 49.2 53.2 33.9 31.9 57.6 55.3 78.0 68.1
Information Querying 66.8 65.9 58.9 55.6 48.3 50.0 69.4 68.3 75.1 72.2
Sports 49.6 50.5 54.6 47.6 45.4 45.6 64.7 66.0 63.9 62.1
Gaming 63.1 61.2 53.4 49.4 41.7 43.5 58.3 64.7 63.1 62.4
Shopping 72.2 67.6 62.7 65.8 51.6 54.1 74.6 76.6 82.5 81.1
Audio Type Natural Sound 57.4 64.6 47.1 50.8 29.4 32.3 54.4 61.5 64.7 61.5
Speech 63.5 61.4 58.7 55.2 51.6 55.6 69.3 70.3 76.2 73.1
Music 69.2 72.0 53.8 56.0 7.7 24.0 69.2 56.0 76.9 76.0
Artificial Sound 60.0 74.0 36.0 40.0 20.0 26.0 56.0 58.0 56.0 68.0
Mixed Sounds 65.3 61.5 56.5 55.5 44.7 45.4 66.0 63.8 69.1 67.9
Topics Humanities and Society 70.7 71.3 63.9 65.9 48.9 51.9 77.4 79.1 82.7 80.6
Geography and Travel 68.8 68.8 56.2 56.2 62.5 68.8 50.0 68.8 75.0 75.0
Technology and Gaming 64.9 64.0 52.7 49.1 42.0 41.2 61.1 62.3 65.6 64.9
Science and Knowledge 57.1 63.0 39.3 51.9 39.3 37.0 53.6 63.0 67.9 70.4
Movies, TV and Animations 61.2 61.2 35.8 37.3 29.9 44.8 55.2 59.7 59.7 65.7
Arts and Performance 58.6 60.7 48.3 42.9 34.5 39.3 58.6 57.1 62.1 71.4
Pets and Animals 64.4 66.7 55.6 48.9 40.0 44.4 64.4 60.0 71.1 73.3
Business and Commerce 73.3 68.6 68.6 71.4 55.8 57.1 79.1 77.1 86.0 87.1
Life and Practical Skills 72.1 64.7 69.2 60.0 54.8 58.8 77.9 64.7 84.6 77.9
Sports and Adventure 53.2 56.9 52.4 48.6 43.7 45.9 65.1 69.7 63.5 63.3
Social Trends and Reactions 51.9 50.0 40.7 46.2 38.9 48.1 57.4 57.7 59.3 55.8
Vehicles and Transportation 55.6 55.6 55.6 44.4 22.2 11.1 55.6 44.4 66.7 55.6
Low Quality and Extended Content 41.7 41.7 54.2 54.2 29.2 37.5 37.5 50.0 66.7 62.5
Multimodal Abilities Visual Reasoning 63.2 65.7 54.8 57.4 43.4 46.6 64.5 64.2 72.8 71.1
Emotional Inference 67.4 68.3 55.8 51.2 37.2 34.1 62.8 58.5 72.1 73.2
Situational Reasoning 70.0 77.8 55.0 50.0 50.0 44.4 70.0 66.7 80.0 77.8
Causal Reasoning 65.7 60.3 47.8 43.1 37.3 48.3 67.2 65.5 70.1 60.3
Factual Recall 64.3 61.0 57.9 55.1 46.2 48.6 68.2 69.0 72.5 71.7
Spatial Reasoning 37.9 45.8 51.7 33.3 41.4 41.7 58.6 45.8 55.2 50.0
Relationship 62.5 62.5 43.8 62.5 50.0 75.0 62.5 75.0 62.5 75.0
Temporal Distance 66.7 81.8 58.3 54.5 58.3 72.7 58.3 81.8 83.3 81.8
Overall 63.5 62.8 55.6 53.9 44.5 47.9 66.3 66.5 71.9 70.5
Table 2: Performance comparison of Gemini 1.5 Pro across different modalities, difficulties, and categories.
We select a wide range of open-sourced and proprietary MLLMs on the MAVERIX benchmark. We report the quantitative results grouped by difficulties in Tab. 1 for their performances with and without subtitles and by different video lengths. The detailed quantitative GPT consistency scores can by found in Tab. D.2, and visualized in Fig. 6. We then showcase the fine-grained quantitative evaluations of Gemini 1.5 Pro in different multimodal input settings and grouped by the splits in Tab. 2. Refer to Section D for additional experimental results. Below, we provide our key observations.
Figure 6: Consistency of model responses on open-ended questions, measuring how reliably models provide coherent and relevant answers across different inputs.
[…]
Are Audio and Text Equally Informative? Refer to Tab. 2 for detailed results. We evaluated Gemini 1.5 Pro, which was selected for its support for audio, text, and video modalities. Across all categories, Gemini 1.5 Pro consistently achieves higher scores with audio-only input compared to text-only input, suggesting that the model may extract more informative or discriminative cues from audio signals than from textual descriptions in this setting. Furthermore, when comparing multimodal configurations, video+audio yields the highest overall performance, surpassing video+subtitle.
Although video + audio generally provide the most significant benefit, certain categories, including sentiment understanding, situational reasoning, causal reasoning, and egocentric agent tasks, experience particularly large performance gains under this configuration. These tasks likely benefit from the richer contextual and temporal cues present in audio, such as tone, ambient sounds, and speaker intent, which are difficult to fully capture through visual or textual modalities alone.
Misaligned Category Trends Between MLLMs and Humans: The category-level analysis in Fig. 1 reveals notable differences in how MLLMs and humans perform across tasks. The quantitative comparisons are detailed in Tab. D.6. Humans tend to excel in dynamic contextual categories such as egocentric reasoning, sports analysis, and gaming strategy, which require interpreting spatial cues, tracking actions over time, and understanding complex interactions. However, their performance is less consistent in domains like shopping and social interaction. Shopping tasks may prioritize factual or procedural details that are less engaging or familiar to humans. For social interaction tasks, inherent subjectivity, such as interpreting cultural contexts, intentions, or unspoken tensions, can lead to diverse and sometimes conflicting human judgments, reducing overall precision.
MLLMs show a different trend. They perform comparatively well on shopping-related tasks, which often involve concrete retrieval of explicit audiovisual details like product descriptions, prices, or transactional cues. These structured object-centric queries align well with the strengths of large models trained on web-scale corpora rich in product and factual information. In contrast, MLLMs consistently struggle in categories that require nuanced temporal reasoning, social understanding, and dynamic context integration, such as predicting team strategies in games or recognizing subtle emotional cues in conversation.
These trends suggest that current MLLMs are more adept at tasks grounded in static, factual associations but remain limited in handling complex, situational reasoning that requires modeling interactions, intent, and evolving context. Closing this gap will likely require advances in temporal modeling, sentiment cue understanding, and the ability to generalize beyond explicit audiovisual patterns.
4 Related Work
MLLM Benchmarks: The development of MLLM benchmarks has progressed from static image-based question answering (QA) to dynamic video understanding. Early benchmarks for evaluating vision-language models (VLMs) focused primarily on images, emphasizing visual grounding and object recognition. Earlier datasets  [11, 3, 59, 41] provided human-annotated captions and QA pairs to assess these core capabilities. Building on this foundation, subsequent benchmarks [42, 34, 34, 44] introduced more domain-specific and knowledge-intensive tasks. A-OKVQA [43] challenges abilities to read and interpret texts within images. More recently, MMMU [54] and MMMU-Pro [55] have expanded these efforts with diverse tasks designed to probe a model’s ability to simultaneously ‘see” and read” from prompted images.
Despite their advances, image-based benchmarks lack the temporal and acoustic context necessary for richer multimodal understanding. This limitation motivated the shift toward video-based evaluation. Fine-grained temporal tasks assessing a model’s ability to perceive motion and temporal sequences have been included in video benchmarks [29, 16, 40, 25, 22, 45, 35, 30]. However, these datasets and recent extensions  [37, 12, 21, 36, 46] tend to feature short clips with constrained domains and limited semantic diversity. They primarily probe local temporal, visual semantics, and factual information, offering limited insight into how models handle the complexities of daily activities, situational awareness, or social reasoning. Most recent works [9, 29, 40, 25, 30, 10] take a step towards everyday contexts by incorporating human actions, but their scopes are restricted to human-centric activities and do not handle visual-audio cues. Video-MME [17] is a comprehensive video benchmark with short, medium, and long videos. However, it only employed the MCQ format, overlooking the quality and richness of free-form expressions generated by MLLMs from open-ended questions.
MAVERIX extends prior benchmarks by explicitly elevating audio as a primary modality, with its questions requiring complex cross-modal reasoning, such as resolving conflicts between dialogue and visual cues. Unlike existing datasets, it systematically evaluates a model’s ability to integrate complete multimodal signals through both descriptive open-ended answers as well as multiple-choice questions, comparisons shown in Tab. 3.
Video Understanding Models: Models based on contrastive pre-training over image-text pairs have demonstrated significant efficacy in learning transferable representations across various visual tasks. These methods have been adapted to specialized domains, including multi-disciplinary scientific understanding and multimodal reasoning. Building upon advancements in pre-training, subsequent work [50, 6, 48, 56, 57, 32] has sought to enhance the multimodal reasoning capabilities of complex spatio-temporal video data and improve their alignments with human preferences through post-training reinforcement learning from human feedback (RLHF). Other efforts [51, 31, 15, 47, 7] integrating a mixture of experts (MoE) architectures [33], which conditionally activates subsets of specialized “expert” networks during processing, have shown improved scalability and performance without a proportional increase in computational costs. However, whether these developments enable models to reason in ways comparable to or surpass humans remains an open question [1, 4, 8, 24], as the assessment of human-like multimodal reasoning, particularly in real-world visual-audio contexts, remains underexplored.
[…]
Benchmarks #Videos Len. #Questions Subtitle Audio Open Multi-Split
MSRVTT-QA 2,990 15.2 72,821
MSVD-QA 504 9.2 50,505
TGIF-QA 9,575 3.6 165,165
ActivityNet-QA 5,800 14.9 58,000
TVQA 2,179 15.2 152,545
How2QA 1,517 11.5 71,812
STAR 914 11.9 59,871
NExT-QA 4,269 12.1 52,631
[…]
Video-Bench 5,917 56.0 17,000
EgoSchema 6,365 18.0 36,054
AutoEval-Video 327 32.1 450
TempCompass 410 11.4 1,540
Video-MME 900 1071.9 2,700
HourVideo 500 2742 12,976
[…]
MAVERIX 700 344.9 2,556
Table 3: Comparison of video QA benchmarks with various attributes.
5 Conclusion
Agentic tasks such as assisting collaborative work and navigating dynamic environments require strong audiovisual reasoning capabilities, and yet these abilities have not been fully assessed in the recent MLLMs. To address this, we introduce MAVERIX, a benchmark designed to evaluate models in complex real-world audiovisual scenarios. It contains 700 videos and 2,533 human carefully crafted questions providing not only a challenging benchmark but also a comprehensive evaluation suite. We conduct a thorough benchmarking of mainstream closed-source and open-source MLLMs on the proposed benchmark. Our evaluations reveal that cross-modal video-audio understanding is essential for solving real-world perception problems. Additionally, they show that even state-of-the-art MLLMs still exhibit a performance gap compared to humans, particularly in socially grounded or dynamic tasks. We hope MAVERIX will serve as a foundation for advancing research in multimodal reasoning and guiding future work towards more context-aware and socially intelligent systems.
Supplementary Material
MAVERIX Supplementary Material
  • Section A: MAVERIX Release v1.0
  • Section B: Rationale and Liminations
  • Section C: Qualitative Comparisons by Modalities
  • Section D: Additional Experiments
    • Section D.1: Standardized Prompt Design
    • Section D.2: Performance by Difficulties
    • Section D.4: Quantitative GPT Evaluations
    • Section D.5: Performance by Categories
    • Section D.6: Common Error Modes
  • Section E: Details on the Human Survey
  • Section F: Additional Data Distribution Analysis
Appendix A MAVERIX Release v1.0
We are releasing MAVERIX v1.0, our proposed benchmark dataset for multimodal video-language understanding. The benchmark dataset is provided as a single JSON file for ease of use and straightforward integration with existing benchmarking pipelines. For each video, the dataset includes metadata and contains multiple-choice and open-ended questions covering multiple tasks from our proposed task suite. Each task is accompanied by a set of questions designed to evaluate visual reasoning, situational awareness, and cross-modal understanding. Relevant timestamps are provided to allow precise video trimming.
We provide all 700 video clips used in our benchmark, and users can download the corresponding videos from the Ego4D website after reviewing and accepting the Ego4D license agreement. We also provide sample videos with annotations from MAVERIX. All materials are available through the project page upon acceptance.
Structure: MAVERIX v1.0 release is organized as follows:
  • data/
    • MAVERIX_v1_0.json: Contains all 2556 questions in the benchmark dataset.
    • vid_only/: Contains all muted videos for evaluations without audio modality.
    • vid/: Contains all videos for evaluations without audio modality.
    • subtitles/: Contains all subtitles for evaluations requiring subtitle access.
    • sample_annotations/: Given that MAVERIX is an evaluation benchmark, ground truth annotations are not released to the public. For review purposes, we provide ground truth annotations for select sample videos in an .json file.
  • src/
    • video_utils.py: A script for dataloading, processing, and evaluation functionalities.
    • _default_template_yaml: A file to specify the default processing template for the evaluation. The data loading path, cache directory, and evaluation modes can be specified here.
    • MAVERIXXX_dev: Individual setting files specifying the test settings for different modes. The currently supported modes are: audio_only, video_only, text_only, video_text, and video_text_audio.
    • Remark: Except for the open-sourced models listed in the paper or available through LMMs-Eval[58, 26], all other experiments require access to proprietary models.
Documentation:
  • We provide a comprehensive datasheet explaining the benchmark dataset’s purpose and intended usage.
License:
  • MAVERIX will be made publicly available under the MIT License. Do note that Ego4D videos are publicly available under the Ego4D License.
Versioning and Updates:
  • We will maintain MAVERIX, with all updates and new versions announced publicly.
Appendix B Rationale and limitations
Broader Impact.
Robust audiovisual understanding is critical for the next generation of agentic systems that interact with humans and environments in real-time. Benchmarks like MAVERIX can help guide the development of models capable of supporting complex tasks such as collaborative decision-making, social interaction understanding, embodied navigation, and context-aware assistance. By evaluating models on realistic, multimodal scenarios, MAVERIX encourages progress toward AI systems that can reason beyond static or unimodal inputs, making them better suited for applications like personal assistants, robotics, and human-computer interaction.
However, improved performance on MAVERIX may not directly translate to safe or ethical deployment. Many categories in MAVERIX, such as social sentiment and situational awareness, involve subjective interpretation and context-sensitive judgment. Without careful handling, these capabilities risk reinforcing biases, misunderstanding cultural cues, or misinterpreting social dynamics, especially when applied across diverse real-world contexts. We encourage users of MAVERIX to consider these risks and to accompany performance improvements with a thorough evaluation of fairness, robustness, and societal impact.
Limitations.
While MAVERIX advances the evaluation of multimodal reasoning, it has several limitations. First, the benchmark focuses on short- to medium-length video segments, which may not fully capture the long-horizon dependencies found in extended real-world interactions. Second, although MAVERIX emphasizes modality interdependence, complete isolation of modalities is challenging, and some tasks may still be partially solvable through strong unimodal priors. Third, certain categories, such as social interaction, show lower human agreement due to the inherent subjectivity of emotional and interpersonal interpretation, which may limit the upper bound of achievable accuracy. Finally, while MAVERIX supports open-ended evaluation using GPT-4o as an automatic judge, automated scoring of generative responses remains imperfect and may introduce noise or bias in assessments.
Despite these limitations, MAVERIX offers a step toward a more comprehensive and realistic evaluation of multimodal LLMs, with the goal of driving future work in building models capable of deeper, contextually grounded, and socially aware reasoning.
Appendix C Qualitative Comparisons by Modalities
We conduct a comprehensive analysis of the performance of Gemini 1.5 Pro, providing a breakdown across each custom-defined split and input modality to better understand its strengths and limitations throughout the benchmark. Specifically, we evaluate the model under different input configurations like Audio Only, Video Only, and Audio+++Video, as well as across diverse content-based splits covering topics, agentic task types, audio characteristics, and difficulty levels. The corresponding results are shown in Fig. C.1 for modality comparison, Fig. C.2 for topic-wise performance, Fig. C.3 for agentic task types, Fig. C.4 for audio type variation, and Fig. C.5 for difficulty levels. This detailed analysis highlights how well Gemini 1.5 Pro generalizes across different challenge dimensions and input conditions.
Figure C.1: Evaluation of multimodal abilities by Categories.
Figure C.2: Evaluation of multimodal abilities by Topics.
Figure C.3: Evaluation of multimodal abilities by Agentic Tasks.
Figure C.4: Evaluation of multimodal abilities by Audio Types.
Figure C.5: Evaluation of multimodal abilities by the level of difficulties.
Appendix D Additional experiments
This section presents extended experimental results to further analyze model performance across various dimensions of the benchmark. In addition to model evaluation, we provide human performance baselines segmented by difficulty levels, input modalities, and agentic task categories for direct comparison.
D.1 Standardized Prompt Design
To ensure fairness in evaluation, we adopt a standardized prompt format across all models. This eliminates potential biases introduced by prompt engineering and ensures that differences in performance are attributed to model capabilities rather than variations in input phrasing. For tasks requiring multiple modalities, the prompts incorporate the modalities in the inputs. In multiple-choice settings, the prompt is presented as follows: {mdframed} [images] + [audio] (if available) This video’s subtitles are listed below: OR No subtitles available. [Subtitles] Select the best answer to the following multiple-choice question based on the video and the subtitles. Respond with only the letter (A, B, C, D, E, F, G, or H) of the correct option. Respond with only the letter (A, B, C, or D) of the correct option. [Question] The best answer is:
For open-ended questions, the following prompt is used:
{mdframed}
[images] + [audio] (if available) This video’s subtitles are listed below: OR No subtitles available. [Subtitles] Select the best answer to the following open-ended question based on the video and the subtitles. [Question] The best answer is:
D.2 Model Performance by Difficulties
We evaluate model performance across different difficulty levels to assess robustness and generalization from simpler to more complex scenarios. The results are presented in Figure D.1, which illustrates how the model’s accuracy declines as task complexity increases and showcases its relative performance compared to human participants.
Figure D.1: Probed model performance across different difficulty levels. This analysis illustrates how the models’ accuracy changes with increasing task complexity and their relative positions against humans.
D.3 Human Performance by Categories
We report performance across distinct agentic task categories to showcase the strengths and weaknesses of the human study participants within the defined contexts. We note that these statistics may be biased due to the composition of the participant pool, and further details on the recruitment process can be found in Sec. E.
Modality Task Worst (%) Best (%) Average (%)
Visual Social 46.15 66.67 54.87
Sports 20.00 100.00 51.43
Information Querying 53.85 71.43 62.15
Sentiment 20.00 73.33 47.71
Shopping 35.71 75.00 57.68
Gaming 37.50 100.00 58.24
Overall 35.20 81.07 55.68
Audio Social 16.67 40.00 24.17
Sports 20.00 42.86 33.57
Information Querying 23.08 36.67 33.27
Sentiment 21.43 40.00 35.36
Shopping 23.08 60.00 44.10
Gaming 25.00 25.00 25.00
Overall 21.88 40.42 32.58
Both Social 66.67 100.00 79.37
Sports 42.86 100.00 80.95
Information Querying 73.33 89.66 81.00
Sentiment 86.67 93.33 89.17
Shopping 60.00 100.00 80.00
Gaming 57.14 87.50 71.13
Overall 64.78 95.08 80.94
Table D.1: Worst, Best, and Average human performance (%) across tasks and modalities.
D.4 Quantitative GPT Evaluations
To further assess model performance beyond multiple-choice accuracy, we conduct a quantitative evaluation of the open-ended question responses using GPT-based scoring. Specifically, as discussed in Sec. 2.2, we prompt GPT-4o to act as an automated grader, evaluating the correctness and relevance of the model’s free-form answers based on predefined guidelines consistent with the task objectives.
This automatic assessment provides an additional perspective on the model’s generative capabilities, capturing nuances that are not reflected in standard accuracy metrics. The results offer insight into how well the model can produce contextually appropriate and informative responses when not constrained by fixed answer options.
We report the GPT-assigned scores across the full benchmark as well as per agentic task category, enabling a more granular understanding of the model’s strengths and weaknesses in open-ended scenarios. These findings complement the multiple-choice evaluations and expose areas where the model demonstrates strong language generation and areas where further improvement is needed. However, we note that these scores may be biased due to limitations in GPT-4o’s own reasoning and evaluation capabilities [28], which can introduce alignment artifacts or systematic preferences.
Models Overall
Human (Video) 1.03
Human (Audio) 0.65
Human (Video+Audio) 2.50
w/o subs w/ subs
Open source MLLMs w/ Truncated Videos
InternVL2 0.896 1.055
Qwen2.5-VL 1.651 1.929
LLaVA-OneVision 1.774 1.124
DeepSeek-VL2-Tiny 0.839 0.870
DeepSeek-VL2-Small 1.128 1.471
Closed-source MLLMs w/ Truncated Videos
GPT-4o 1.327 2.308
o1 2.697 3.075
Claude 3.5 Sonnet 1.964 2.568
Grok2-Vision 2.193 2.765
Nova-Lite 1.444 1.859
Nova-Pro 1.998 2.364
Gemini 1.5 Pro 2.170 2.538
Open source MLLMs w/ Full-Length Videos
InternVL2 0.895 1.102
Qwen2.5-VL 1.489 1.830
LLaVA-OneVision 1.471 1.666
DeepSeek-VL2-Tiny 0.675 0.808
DeepSeek-VL2-Small 0.839 1.295
Closed-source MLLMs w/ Full-Length Videos
GPT-4o 1.061 2.148
o1 2.484 2.983
Claude 3.5 Sonnet 0.017 0.401
Grok2-Vision 1.935 2.622
Nova-Lite 1.653 1.708
Nova-Pro 1.615 2.012
Gemini 1.5 Pro 1.956 2.318
Table D.2: GPT-4o evaluation of consistency score.
Models Overall
Human (Video) 1.02
Human (Audio) 0.58
Human (Video+Audio) 2.71
w/o subs w/ subs
Open source MLLMs w/ Truncated Videos
InternVL2 0.129 1.007
Qwen2.5-VL 1.404 1.702
LLaVA-OneVision 1.168 1.005
DeepSeek-VL2-Tiny 0.562 0.614
DeepSeek-VL2-Small 1.087 1.851
Closed-source MLLMs w/ Truncated Videos
GPT-4o 1.277 2.103
o1 2.484 2.803
Claude 3.5 Sonnet 2.122 2.283
Grok2-Vision 1.901 2.434
Nova-Lite 1.287 1.602
Nova-Pro 1.518 1.741
Gemini 1.5 Pro 1.970 2.299
Open source MLLMs w/ Full-Length Videos
InternVL2 1.333 0.975
Qwen2.5-VL 1.275 1.660
LLaVA-OneVision 1.277 1.572
DeepSeek-VL2-Tiny 0.687 1.168
DeepSeek-VL2-Small 0.873 1.268
Closed-source MLLMs w/ Full-Length Videos
GPT-4o 1.277 2.103
o1 2.484 2.803
Claude 3.5 Sonnet 0.319 0.415
Grok2-Vision 1.693 2.189
Nova-Lite 1.196 1.481
Nova-Pro 1.384 1.678
Gemini 1.5 Pro 1.808 2.136
Table D.3: GPT-4o evaluation of correctness score.
Models Overall
Human (Video) 0.90
Human (Audio) 0.52
Human (Video+Audio) 2.55
w/o subs w/ subs
Open source MLLMs w/ Truncated Videos
InternVL2 0.586 1.120
Qwen2.5-VL 1.746 2.033
LLaVA-OneVision 0.953 1.090
DeepSeek-VL2-Tiny 0.653 0.68
DeepSeek-VL2-Small 1.162 1.493
Closed-source MLLMs w/ Truncated Videos
GPT-4o 1.284 2.220
o1 3.012 3.236
Claude 3.5 Sonnet 1.984 2.613
Grok2-Vision 2.454 2.968
Nova-Lite 1.339 1.701
Nova-Pro 1.812 1.841
Gemini 1.5 Pro 2.374 2.716
Open source MLLMs w/ Full-Length Videos
InternVL2 1.107 1.159
Qwen2.5-VL 1.559 1.996
LLaVA-OneVision 1.40 1.679
DeepSeek-VL2-Tiny 0.839 0.953
DeepSeek-VL2-Small 0.970 1.382
Closed-source MLLMs w/ Full-Length Videos
GPT-4o 1.032 1.967
o1 2.880 3.159
Claude 3.5 Sonnet 0.146 0.494
Grok2-Vision 2.237 2.792
Nova-Lite 1.273 1.562
Nova-Pro 1.567 1.832
Gemini 1.5 Pro 2.288 2.618
Table D.4: GPT-4o evaluation of detailed orientation score.
Models Overall
Human (Video) 1.26
Human (Audio) 0.58
Human (Video+Audio) 2.71
w/o subs w/ subs
Open source MLLMs w/ Truncated Videos
InternVL2 0.129 1.393
Qwen2.5-VL 2.007 2.249
LLaVA-OneVision 1.168 1.311
DeepSeek-VL2-Tiny 0.839 0.870
DeepSeek-VL2-Small 1.505 1.851
Closed-source MLLMs w/ Truncated Videos
GPT-4o 1.586 2.499
o1 3.032 3.291
Claude 3.5 Sonnet 2.122 2.734
Grok2-Vision 2.468 2.960
Nova-Lite 1.742 2.039
Nova-Pro 2.046 2.238
Gemini 1.5 Pro 2.435 2.801
Open source MLLMs w/ Full-Length Videos
InternVL2 1.332 1.366
Qwen2.5-VL 1.818 2.187
LLaVA-OneVision 1.725 1.995
DeepSeek-VL2-Tiny 1.067 1.168
DeepSeek-VL2-Small 1.292 1.655
Closed-source MLLMs w/ Full-Length Videos
GPT-4o 1.274 2.250
o1 2.912 3.214
Claude 3.5 Sonnet 0.319 0.533
Grok2-Vision 2.240 2.751
Nova-Lite 1.653 1.936
Nova-Pro 1.915 2.2348
Gemini 1.5 Pro 2.257 2.618
Table D.5: GPT-4o evaluation of context score.
D.5 Quantitative Evaluations by Categories
To provide a comprehensive understanding of model performance across the diverse challenge dimensions of the benchmark, we report quantitative results for all probed models across each defined split. This includes evaluations broken down by agentic task categories, difficulty levels, audio characteristics, topics, and input modalities.
These category-wise results allow us to examine how different models handle specific functional scenarios, content types, and levels of complexity. Through these experiments, we hope to reveal both general patterns and areas of strength or weakness. By comparing performance across these splits, we observe how well models generalize to varying real-world situations and identify which aspects of the benchmark remain most challenging.
Together, these fine-grained evaluations provide a detailed view of model behavior across the full spectrum of tasks and shed light on future work on multimodal understanding and generalization.
GPT-4o o1 Gemini 1.5 Pro Nova Pro Grok2 Claude 3.5 Sonnet
Difficulty Easy 69.8 79.3 69.5 51.5 65.1 68.9
Medium 67.0 76.0 69.4 47.5 61.2 64.6
Hard 49.6 67.4 49.6 34.1 47.4 44.4
Agentic Tasks Social 64.6 69.7 62.6 49.5 60.6 66.7
Sentiment 56.8 74.1 66.7 46.9 51.9 51.9
Egocentric Agent 62.7 69.5 57.6 40.7 54.2 64.4
Information Querying 67.5 75.5 69.4 47.2 62.6 64.9
Sports 66.4 79.0 64.7 40.3 58.8 56.3
Gaming 64.1 72.8 58.3 54.4 52.4 59.2
Shopping 68.3 85.7 74.6 47.6 73.0 73.0
Audio Type Natural Sound 57.4 66.2 54.4 35.3 50.0 48.5
Speech 67.7 76.7 69.3 44.8 61.7 65.0
Music 57.7 61.5 69.2 38.5 57.7 57.7
Artificial Sound 70.0 78.0 56.0 58.0 54.0 60.0
Mixed Sounds 63.4 78.2 66.0 52.3 63.0 64.9
Topics Humanities and Society 71.4 78.9 77.4 51.1 66.9 72.2
Geography and Travel 56.2 68.8 50.0 37.5 43.8 43.8
Technology and Gaming 65.6 71.8 61.1 50.4 53.4 57.3
Science and Knowledge 60.7 64.3 53.6 28.6 50.0 53.6
Movies, TV and Animations 55.2 68.7 55.2 40.3 43.3 55.2
Arts and Performance 51.7 65.5 58.6 41.4 58.6 75.9
Pets and Animals 62.2 68.9 64.4 48.9 62.2 57.8
Business and Commerce 72.1 86.0 79.1 60.5 76.7 75.6
Life and Practical Skills 74.0 79.8 77.9 51.9 71.2 69.2
Sports and Adventure 63.5 80.2 65.1 46.0 57.1 57.9
Social Trends and Reactions 55.6 77.8 57.4 33.3 61.1 59.3
Vehicles and Transportation 88.9 100.0 55.6 55.6 55.6 55.6
Low Quality and Extended Content 54.2 58.3 37.5 16.7 50.0 54.2
Multimodal Abilities Visual Reasoning 59.2 73.2 64.5 42.1 57.5 59.6
Emotional Inference 60.5 69.8 62.8 46.5 53.5 48.8
Situational Reasoning 65.0 80.0 70.0 40.0 75.0 75.0
Causal Reasoning 65.7 73.1 67.2 50.7 58.2 68.7
Factual Recall 68.9 79.2 68.2 49.2 63.2 65.0
Spatial Reasoning 62.1 62.1 58.6 41.4 44.8 55.2
Relationship 75.0 81.2 62.5 50.0 81.2 75.0
Temporal Distance 66.7 66.7 58.3 58.3 50.0 66.7
Overall 65.4 75.9 66.3 46.9 60.6 63.1
Table D.6: Performance comparison across different models and categories on truncated videos.
GPT-4o o1 Gemini 1.5 Pro Nova Pro Grok2 Claude 3.5 Sonnet
Difficulty Easy 64.8 73.0 69.4 58.2 59.2 17.4
Medium 62.2 72.4 68.0 58.1 55.2 14.8
Hard 46.5 65.4 55.1 41.7 53.5 17.3
Agentic Tasks Social 66.7 65.6 59.4 63.5 62.5 9.4
Sentiment 54.3 67.9 64.2 46.9 55.6 24.7
Egocentric Agent 55.3 61.7 55.3 19.1 34.0 14.9
Information Querying 63.9 75.4 68.3 60.7 57.9 17.5
Sports 55.3 69.9 66.0 55.3 59.2 14.6
Gaming 60.0 71.8 64.7 48.2 48.2 22.4
Shopping 60.4 75.7 76.6 64.0 62.2 10.8
Audio Type Natural Sound 70.5 79.8 61.5 53.8 43.1 30.8
Speech 56.2 81.2 70.3 57.3 60.7 14.4
Music 55.3 70.2 56.0 48.0 56.0 16.0
Artificial Sound 51.9 59.3 58.0 48.0 48.0 28.0
Mixed Sounds 58.2 59.7 63.8 55.0 54.6 12.8
Topics Humanities and Society 57.1 67.9 79.1 66.7 64.3 27.1
Geography and Travel 60.0 68.9 68.8 56.2 56.2 18.8
Technology and Gaming 61.4 72.9 62.3 50.9 50.0 17.5
Science and Knowledge 69.4 75.3 63.0 55.6 51.9 11.1
Movies, TV and Animations 55.0 68.8 59.7 52.2 46.3 11.9
Arts and Performance 57.7 75.0 57.1 46.4 57.1 17.9
Pets and Animals 44.4 88.9 60.0 57.8 57.8 17.8
Business and Commerce 62.5 62.5 77.1 65.7 67.1 12.9
Life and Practical Skills 57.8 68.6 64.7 57.6 62.4 8.2
Sports and Adventure 43.9 65.9 69.7 52.3 55.0 16.5
Social Trends and Reactions 72.2 77.8 57.7 51.9 51.9 15.4
Vehicles and Transportation 53.4 63.8 44.4 44.4 33.3 0.0
Low Quality and Extended Content 63.3 73.9 50.0 20.8 50.0 8.3
Multimodal Abilities Visual Reasoning 62.5 66.7 64.2 51.5 54.4 14.7
Emotional Inference 68.8 75.0 58.5 53.7 43.9 19.5
Situational Reasoning 81.8 90.9 66.7 55.6 77.8 16.7
Causal Reasoning 53.8 56.9 65.5 55.2 55.2 25.9
Factual Recall 63.8 71.9 69.0 58.1 58.8 15.9
Spatial Reasoning 52.0 60.0 45.8 41.7 41.7 16.7
Relationship 66.0 82.0 75.0 68.8 68.8 0.0
Temporal Distance 56.4 73.9 81.8 54.5 45.5 18.2
Overall 60.6 71.5 66.5 55.5 56.5 16.3
Table D.7: Performance comparison across different models and categories on full-length videos.
D.6 Common Error Modes
To better understand the limitations of the evaluated models, we analyze common error patterns observed across the benchmark. We find that model failures often arise from modality-specific weaknesses, such as misinterpretation of complex visual cues, inability to process nuanced audio signals, or challenges in aligning multimodal information with nuanced context. Common failure modes are showcased in Fig. D.2,  D.3,  D.4,
Figure D.2: Error analysis showing that o1 fails to correctly answer the question when the audio cannot be transcribed into text-based subtitles, leading to an incorrect connection between the cityscape video and the instrumental music.
Figure D.3: Error analysis showing inconsistent responses from o1, where it failed to accurately identify Hika-chan, when the question was rephrased slightly.
Figure D.4: Error analysis highlighting o1’s failure in correctly synchronizing the audio with visual cues, leading to misinterpretation of the woman picking up tree ornaments.
Figure D.5: Error analysis highlighting o1’s contextual misinterpretation, attributing the man’s smile to a general amusing memory rather than the dog’s barking resembling gunfire. Additionally, Gemini Pro 1.5 exhibits visually induced audio hallucination, mistaking the dog’s open-mouth appearance for barking and associating it with a crying baby.
Figure D.6: Error analysis showing o1 and Gemini Pro 1.5 both struggling with spatial misjudgment and temporal misalignment, leading to errors in estimating the number of defenders due to difficulties in distinguishing actual close defenders and potential misalignment with the described action.
Figure D.7: Error analysis showing Gemini Pro 1.5 providing inconsistent responses despite clear visuals and audio cues, failing to accurately identify the cause of the glass breaking due to rephrased question structure.
Appendix E Questionnaire Survey
To establish human performance baselines for comparison with the probed models, we conducted a questionnaire study approved by our institutional IRB. We recruited participants with proficiency in at least one language present in the dataset and prior familiarity with video content to ensure they could reasonably complete the tasks. Participants answered questions presented in the same format as those used for model evaluation.
The survey included a total of 200 questions curated from the dataset, covering diverse subcategories within each defined split. For each participant, the questionnaire consisted of two parts: one multiple-choice question (MCQ) with eight answer options, followed by one open-ended question from a different video. To minimize content overlap and reduce potential bias, no video was reused between questions within a single questionnaire. This approach ensured broad coverage of the dataset’s taxonomy while avoiding contamination across tasks.
For evaluation, MCQ responses were scored against the ground truth to calculate accuracy, while open-ended answers were assessed using the same GPT-4o evaluation pipeline applied to model outputs, ensuring consistency across human and model scoring.
We tested human performance across three different conditions: audio only, visual only, and audio plus video. Each condition was toggle-enabled within the survey toolkit, allowing participants to be assigned to a specific modality. The survey interfaces for the conditions are shown in Fig. E.2 for audio only, Fig. E.3 for visual only, Fig. E.1 for both modalities. To prevent cross-condition contamination and ensure focused evaluation, each participant completed the entire study under only one selected modality.
This human study provides us with valuable reference points for interpreting model results and understanding how humans perform under comparable multimodal constraints.
Figure E.1: Screenshot of the questionnaire under the audio-plus-visual condition, showcasing the interface used to assess human performance across multiple-choice questions. Results offer a baseline for comparison with multimodal model performance across diverse tasks and difficulty levels.
Figure E.2: Questionnaire screenshot showing open-ended questions under the audio-only condition. Responses were evaluated using the GPT-4o grading pipeline, capturing the quality of language generation based solely on audio context.
Figure E.3: Questionnaire screenshot showing open-ended questions under the visual-only condition. GPT-4o evaluation scores assess human understanding based solely on visual information, without the aid of audio context.
Appendix F Dataset Distribution Continued
In this section, we provide a detailed breakdown of the taxonomy distributions within each defined agentic task category. For each task type, we analyze the composition of its associated taxonomies to illustrate the diversity and balance of the dataset across different functional scenarios. The corresponding distributions are visualized in Fig. F.1 for Gaming, Fig. F.2 for Egocentric Agent, Fig. F.3 for Information Querying, Fig. F.4 for Sentiments, Fig. F.5 for Shopping, Fig. F.6 for Social, and Fig. F.7 for Sports. These visualizations offer insight into the internal structure of each agentic task and highlight the variety of situations represented within the dataset.
Figure F.1: Taxonomy composition in Gaming.
Figure F.2: Taxonomy composition in Egocentric Agent.
Figure F.3: Taxonomy composition in Information Querying.
Figure F.4: Taxonomy composition in Sentiments.
Figure F.5: Taxonomy composition in Shopping.
Figure F.6: Taxonomy composition in Social.
Figure F.7: Taxonomy composition in Sports.

### v2

@@
-MAVERIX: Multimodal Audio-Visual Evaluation Reasoning IndeX
+Liuyue Xie∗1  Avik Kuthiala∗1  George Z. Wei∗1 Ce Zheng1  Ananya Bal1  Mosam Dabhi1  Liting Wen1 Taru Rustagi1  Ethan Lai1  Sushil Khyalia1  Rohan Choudhury1
 Abstract
-Frontier models have either been language-only or have primarily focused on vision and language modalities.
-Although recent advancements in models with vision and audio understanding capabilities have shown substantial progress, the field lacks a standardized evaluation framework for thoroughly assessing their cross-modality perception performance.
-We introduce MAVERIX (Multimodal Audio-Visual Evaluation Reasoning IndeX), a novel benchmark with 700 videos and 2,556 questions explicitly designed to evaluate multimodal models through tasks that necessitate close integration of video and audio information.
-MAVERIX uniquely provides models with audiovisual tasks, closely mimicking the multimodal perceptual experiences available to humans during inference and decision-making processes.
-To our knowledge, MAVERIX is the first benchmark aimed explicitly at assessing comprehensive audiovisual integration.
-Experiments with state-of-the-art models, including Gemini 1.5 Pro and o1, show performance approaching human levels (around 70% accuracy), while human experts reach near-ceiling performance (95.1%).
-With standardized evaluation protocols, a rigorously annotated pipeline, and a public toolkit (maverix-benchmark.github.io) [^1] Project page will be available soon., MAVERIX establishes a challenging testbed for advancing audiovisual multimodal intelligence.
-Figure 1: MAVERIX contains over 2,500 audio-visual tasks spanning diverse domains with tight modality interdependence (left) and remains challenging, as state-of-the-art multimodal frontier models are consistently outperformed by average human participants (right).
-Figure 2: Example questions and answers in the MAVERIX benchmark.
-1 Introduction
+We introduce MAVERIX (Multimodal audiovisual Evaluation and Recognition IndeX), a unified benchmark to probe the video understanding in multimodal LLMs, encompassing video, audio, text inputs with human performance baselines.
+Although recent advancements in models with vision and audio understanding capabilities have shown substantial progress, the field lacks a standardized evaluation framework to thoroughly assess their cross-modality comprehension performance.
+MAVERIX curates 2,556 questions from 700 videos, in the form of both multiple-choice and open-ended formats, explicitly designed to evaluate multimodal models through questions that necessitate tight integration of video and audio information, spanning a broad spectrum of agentic scenarios.
+MAVERIX uniquely provides models with audiovisual questions, closely mimicking the multimodal perceptual experiences available to humans during inference and decision-making processes.
+To our knowledge, MAVERIX is the first benchmark aimed explicitly at assessing comprehensive audiovisual integration in such granularity.
+Experiments with state-of-the-art models, including Qwen 2.5 Omni and Gemini 2.5 Flash-Lite, show performance around 64% accuracy, while human experts reach near-ceiling performance of 92.8%, exposing a substantial gap to human-level comprehension.
+With standardized evaluation protocols, a rigorously annotated pipeline, and a public toolkit, MAVERIX establishes a challenging testbed for advancing audiovisual multimodal intelligence.
+Introduction
 Human cognition seamlessly integrates visual and auditory information to reason, infer, and interact within dynamic environments.
-Replicating this ability in Multimodal Large Language Model (MLLM) systems remains a central challenge for AI, as autonomous agents must process complex audiovisual input to engage meaningfully with the world [32, 4, 1].
+Replicating this ability in Multimodal Large Language Model (MLLM) systems remains a central challenge for AI, as autonomous agents must process complex audiovisual input to engage meaningfully with the world (Lin et al.
+2023; Amirizaniani et al.
+2024; tes 2024).
+Figure 1: An illustration of our proposed benchmark, which includes highly audiovisual correlated questions and paraphrased questions, can be used to evaluate the model’s underlying comprehension abilities and their gaps to humans.
+Figure 2: Example Agentic Categories and corresponding QAs in the MAVERIX benchmark.
 Recent progress in multimodal foundation models has brought us closer to this goal, but current benchmarks fall short in assessing their abilities to reason with multimodal inputs.
-Most benchmarks focus on static images [11, 3], simple recognition, or tasks that can be solved through unimodal cues like transcripts [44, 13].
-These benchmarks fail to probe the deeper, joint reasoning across modalities needed for real-world scenarios, such as interpreting social interactions or anticipating off-screen events [10, 36].
-A core obstacle in designing effective multimodal benchmarks is ensuring that tasks genuinely require multimodal understanding rather than allowing models to exploit unimodal shortcuts.
-Many existing video-language benchmarks reduce to visible-object recognition or dialogue parsing, sidestepping the need to synthesize audio-visual dependencies [40, 25, 29].
-To address this, we introduce MAVERIX, a benchmark designed to evaluate multimodal video-audio understanding through tasks that have tight modality interdependence.
-MAVERIX features four challenging tasks: situational awareness, contextual summarization, social sentiment analysis, and dynamic reasoning, covering 700 videos and over 2,500 carefully designed questions.
-These are constructed through a hybrid human-AI pipeline to ensure that solving them requires intertwined audiovisual synthesis, revealing the underlying multimodal reasoning capabilities of models applied to the benchmark tasks.
-Evaluations of state-of-the-art models, including Gemini 1.5 Pro [49], GPT-4o [39], and o1 [38], reveal significant gaps.
-Gemini 1.5 Pro, even with direct audio-video inputs, achieves 71.9% accuracy on multiple choice questions, marginally lower than human performance (80%).
+Most benchmarks focus on static images (Chen et al.
+2015; Agrawal et al.
+2019; Li et al.
+2025b), simple recognition, or questions that can be solved through unimodal cues, such as transcripts (Singh et al.
+2019; Chen et al.
+2024b).
+These benchmarks fail to probe the deeper, joint reasoning across modalities needed for real-world scenarios, such as interpreting social interactions or anticipating off-screen events (Chandrasegaran et al.
+2024; Mangalam, Akshulakov, and Malik 2023).
+A core obstacle in designing effective multimodal benchmarks is ensuring that the questions genuinely require multimodal understanding rather than allowing models to exploit unimodal shortcuts or common sense from the training data.
+For benchmarks designed to expose the model understanding for highly multimodal data, their common adoption of a 4-way multiple-choice question for evaluation provides limited insights into the underlying interpretations (Li et al.
+2025b; Hong et al.
+2025).
+Many existing video-language benchmarks reduce to visible-object recognition or dialogue parsing, bypassing the need to synthesize audiovisual dependencies (Patraucean et al.
+2023; Kesen et al.
+2023; Li et al.
+2024b).
+To address this, we introduce MAVERIX, a benchmark designed to evaluate multimodal video-audio understanding through questions that have tight modality interdependence.
+MAVERIX features questions from challenging agentic scenario categories: factual recall, causal understanding, sentiment analysis, temporal recall, situational awareness, context understanding, social interaction understanding, and emotional synthesis, covering 700 videos and 2,556 carefully designed questions.
+These are constructed through a hybrid human-AI pipeline to ensure that solving them requires intertwined audiovisual synthesis, revealing the underlying multimodal reasoning capabilities of models applied to the benchmark questions.
+Evaluations of state-of-the-art proprietary and open-source models of different sizes, including Gemini 2.0/2.5 Flash-Lite(FL)(Team and Petko Georgiev 2024), GPT-4o (OpenAI and Aaron Hurst 2024), and Qwen 2.5 Omni (Xu et al.
+2025), reveal significant gaps as shown in Fig.
+1.
+Gemini 2.5 Flash-Lite, even with direct audio-video inputs, achieves 54.7% accuracy on multiple choice questions, significantly lower than human performance (92.8%).
 Open-ended responses further expose weaknesses in temporal reasoning and contextual understanding, with models averaging only 1.9/5 vs.
 human 2.79/5 in GPT-4o-judged scoring.
-Further, models that are not capable of processing raw audio and rely solely on transcripts perform even worse, highlighting the inadequacy of text-only proxies for rich audiovisual comprehension [37, 17].
+Further, models that are not capable of processing raw audio and rely solely on transcripts perform even worse, highlighting the inadequacy of text-only proxies for rich audiovisual comprehension (Ning et al.
+2023; Fu et al.
+2024).
 By providing a unified evaluation framework, high-quality human-validated questions, and an open-source toolkit, MAVERIX aims to advance research toward robust multimodal reasoning at the human level.
-2 Benchmark Design and Constructions
-Designing benchmarks for audiovisual understanding requires tasks that demand genuine cross-modal reasoning, not just unimodal pattern matching.
-In real-world scenarios, one modality often dominates, such as visuals in a silent event or audio in a podcast, but effective understanding depends on integrating cues across modalities.
-Existing benchmarks typically focus on narrow tasks or merge disjointed datasets, limiting their ability to evaluate general-purpose multimodal LLMs (MLLMs).
-Moreover, many video-language benchmarks rely on multiple-choice questions (MCQs) for their simplicity, but small answer spaces of the common four-option MCQs inflate scores through random guessing and fail to assess free-form reasoning — a core aspect of human-like comprehension.
-Figure 3: Annotation workflow outlining the stages of data selection, review, and refinement.
-To address these limitations, MAVERIX introduces a unified benchmark of diverse real-world tasks that systematically test the interdependence of modality while supporting a dual evaluation framework of eight-option MCQs and open-ended questions.
-The expanded eight-option format reduces guess bias (12.5% baseline), providing more reliable accuracy metrics, while open-ended tasks require models to generate nuanced, context-aware responses.
-For open-ended evaluation, we employ GPT-4o as an automated judge, scoring responses against human-curated references using a five-dimensional rubric: correctness of information, attention to detail, contextual understanding, temporal coherence, and consistency under paraphrasing.
-MAVERIX’s taxonomy and revealed performance gaps are shown in Fig.
-1.
-This hybrid design balances scalability and depth: MCQs provide standardized large-scale evaluation, while open-ended questions probe the models’ ability to synthesize and express complex audiovisual reasoning.
-Together, they create a robust assessment ecosystem tailored to intertwined multimodal understanding.
-We detail our pipeline used to generate MAVERIX’s 2,500+ high-quality questions and answers.
-2.1 Dataset Generation Pipeline
-Dataset Collection To ensure broad coverage of multimodal reasoning, we designed four domain hierarchies spanning agentic tasks, topics, video categories, and multimodal abilities, as shown in Fig.
-2.
-Under agentic tasks, we defined seven key areas—social interactions, sentiment expression, information querying, sports analysis, gaming commentary, egocentric navigation, and multilingual contexts—each reflecting real-world scenarios that require complementary audiovisual cues.
-For example, gaming videos involve coordinating visual gameplay with voice chat strategies.
-In total, we annotated 155 fine-grained activity tags (e.g., “referee decision analysis” under sports) to capture diverse reasoning challenges (see Fig.
-5).
-Figure 4: Overview of the dataset distribution. The corresponding number of samples in each category is shown.
-We sourced videos from YouTube-8M [2], MSR-VTT [53], UR-FUNNY-V2 [20], Ego4D [19], and AudioSet [18], balancing short (<1absent1<1< 1 min), medium (2–10 min), and long (10–65 min) formats.
-Each video includes subtitles, categorized audio (e.g., speech, environmental, artificial), and visual keyframes sampled at 1-second intervals.
-After deduplication and quality checks, MAVERIX comprises 700 videos with balanced coverage across domains and durations.
-Multimodal Question-Answer Annotation To evaluate models’ ability to synthesize audiovisual information, we created 852 eight-option multiple-choice questions through a three-stage annotation protocol.
-A team of 11 researchers proficient in English (or another second language) with expertise in vision-language learning ensured high-quality, semantically precise questions.
-The process involved (1) selecting videos and identifying clips requiring understanding multiple modalities, (2) drafting questions designed to be answerable only through joint audiovisual analysis, and (3) applying dual-validation checks to eliminate unimodal shortcuts by verifying that neither audio-only nor vision-only playback sufficed.
-The correct option is manually annotated, then paraphrased and modified into semantically different options with structurally similar alternative options.
-Surviving questions underwent peer review, with ambiguous cases resolved by consensus.
-Figure 5: Dataset statistics.
-Left: distribution of video sources.
-Center: question and answer length across difficulty levels.
-Right: distribution of video durations for full-length and truncated clips.
+Benchmark Design and Construction
+MAVERIX challenges multimodal large language models (MLLMs) to integrate audio and visual evidence under realistic conditions.
+This section elaborates on four key aspects: (i) the motivation behind our design, (ii) the dataset construction pipeline, (iii) the dual-format evaluation protocol, and (iv) dataset statistics.
+Fig.
+3 visualizes the pipeline; Tab.
+1 compares the benchmark components with the relevant works; Tab.
+2 summarizes key dataset statistics.
+Benchmark #Vid.
+Med.
+Len.
+(s) #Q Mod.
+MCQ.
+#Div.
+Diff.
+Shortcut Human OE.
+MSRVTT-QA 2,990 15 72,821 V 4-MCQ 1 MSVD-QA 504 9 50,505 V 4-MCQ 1 ActivityNet-QA 5,800 15 58,000 V 4-MCQ 1 How2QA 1,517 11 71,812 V+S+A 4-MCQ 1 AutoEval-Video 327 32 450 V 4-MCQ 1 TempCompass 410 11 1,540 V 4-MCQ 2 Video-MME 900 1,072 2,700 V+S 4-MCQ 3 OmniBench 1142 I+A 4-MCQ 3 WorldSense 1,662 141.1 3,172 V+S+A 4-MCQ 3 HourVideo 500 2,742 12,976 V+S 4-MCQ 2 MAVERIX 700 106 2,556 V+S+A 8-MCQ 7 MAVERIX-Long 700 345 2,556 V+S+A 8-MCQ 7
+Table 1: Comparison with prior video-question benchmarks.
+Mod.: V (video), S (subtitles), A (audio).
+Ans.: 4-MCQ, “8-MCQ+OE” (eight-option plus open-ended).
+Diff.: crowdsourced or expert difficulty labels.
+#Div.: Number of division types.
+Shortcut: dataset validated against audio-only / video-only ablations.
+Human: ✓ if a benchmark reports any human baseline.
+Design Motivation and Principle
+While previous video-understanding benchmarks curate multiple-choice question-answer pairs over different topics, some still suffer unimodal shortcuts and are limited in exposing models’ underlying biases.
+For evaluating the models’ multimodal understanding abilities, we source videos that capture a wide range of temporal events, spatial motions, and audiovisual correlations.
+While sourcing the videos and constructing the multiple choice and open-ended question-answer pairs, we follow the following design principles.
+Avoid Unimodal Shortcut.
+Many existing image and video based question answering benchmarks (e.g., TVQA, MSR‑VTT Q&A) contain questions that can be answered from captions or a single salient frame, enabling unimodal shortcuts.
+Wide range of evaluation dimensions.
+Deploying MLLMs in the real world requires that the models understand and handle scenarios with different skills.
+Single‑skill benchmarks do not reflect the breadth of reasoning required in open‑world settings.
+The curated benchmark measures the models in six evaluation dimensions, covering acoustic understanding, agentic skills, understanding by broad and sub-topic taxonomy, temporal understanding, and multimodal synthesis abilities.
+Prevention of guess inflation.
+We design a hybrid of eight-way multiple-choice, and open-ended QA for the benchmark, such that by-passing the questions with model-inherent biases can be evidently exposed.
+The hybrid design evaluates the models’ actual abilities to interpret the input sources in different modality settings and provides a fair evaluation of their capabilities.
+Dataset Generation Pipeline
+Video Collection.
+We primarily source our video content from five datasets: YouTube‑8M(Abu-El-Haija et al.
+2016), MSR‑VTT(Xu et al.
+2016), UR‑FUNNY‑V2(Hasan et al.
+2019), Ego4D(Grauman et al.
+2021), and AudioSet(Gemmeke et al.
+2017).
+YouTube‑8M(Abu-El-Haija et al.
+2016) and AudioSet(Gemmeke et al.
+2017) are large-scale datasets covering a wide range of taxonomies, where the videos have strong audiovisual correlations.
+MSR-VTT comprises of high quality video descriptions designed for video translation QAs.
+UR-FUNNY-V2 exhibits videos of different emotional states, challenging the models in their sentimental understanding.
+Lastly, Ego4D consists of egocentric, long-duration videos to probe model’s understanding to daily interactions for agentic scenarios.
+The videos are selected by human annotators according to the principles described above, ensuring a thorough distribution across the topics, with different durations: short with <1<1 min, medium with 1-10 mins, and long with 10-65 mins.
+Each video is processed and accompanied by subtitles generated with Whisper-v3 (Radford et al.
+2022) to ensure a fair evaluation on video-text models without audio-support.
+Initial Question Answering Annotation.
+A team of 8 expert annotators engaged in the initial question-answering pair curation.
+The annotators provided at least one question answer pair to each video to generate the initial ground truths.
+Then the same pair is expanded into eight-way multiple choice question with alternative distractive answers.
+Shortcut Removal and Validation.
+Following the initial annotation, we use a semi-automated approach to validate difficulties with MLLMs and refine the questions to avoid any potential shortcuts.
+Each question undergoes three ablation tests with GPT‑4o-mini and Gemini 2.0-FL: text‑only, video‑only, and videos+subtitles.
+If any ablation yields the correct answer for both models, the item is flagged and revised to reduce reliance on unimodal cues.
+All revisions are logged, and the final set is approved after a second expert pass.
 For example, a valid question might ask, “Why did the mechanic abruptly stop speaking?” requiring both visual cues (e.g., discovering a leak) and audio cues (e.g., sudden silence).
 This protocol ensures MAVERIX’s QA pairs demand genuine modality interdependence, preventing reliance on any single modality.
-Quality Assurance To ensure the reliability of MAVERIX’s videos and annotations, we apply a two-stage quality control protocol combining expert review and model-based validation.
-Each QA pair undergoes four checks by a separate annotator, as illustrated in Fig.
-3: (1) linguistic validity for clear and grammatical phrasing, (2) answerability of whether the question is resolvable via the video’s audiovisual content), (3) option integrity to ensure one correct answer with plausible distractors like semantically tangent or structurally identical options, and (4) modality interdependence, using cross-modal invalidation tests from Section 3.1 (e.g., disabling audio or video to detect shortcuts).
+The difficulty labels were crowd-sourced through Amazon MTurk service (mtu 2005) with 219 unique participants for gauging common consensus, and are determined based on the subtlety of cross-modal cues, the depth of understanding required, and the ease of locating relevant information in the video.
+The human performance evaluations were gathered through MTurk with 382 participants answering a 1/3 subset of the MCQs and open-ended questions.
+Statistic Audio Type Agentic Categories Topic Domain Overall
+Sub-class of QA
+Mixed
+Sound
+Speech
+Artificial
+Sound
+Information
+Querying
+Egocentric
+Agent
+Sentiment
+Analysis
+Humanity &
+Society
+Business &
+Commerce
+Agentic Abilities
+Causal Relationship 66 111 15 75 27 33 27 21 201
+Emotional Inference 51 57 18 66 21 27 27 12 129
+Factual Recall 516 690 60 672 171 186 210 120 1311
+Situational Understanding 27 27 10 33 9 12 12 6 70
+Context Understanding 309 414 45 237 147 138 111 84 771
+QA Lengths
+Question 11.26 11.63 9.06 10.31 10.41 10.00 10.09 10.51 11.28
+Options 11.76 10.52 11.99 9.92 12.29 12.89 10.30 12.54 11.13
+Open-ended answer 13.30 12.69 11.68 10.46 12.77 13.51 12.73 12.12 12.85
+Subtitle length 440.79 682.24 419.87 485.08 1257.19 360.40 351.32 488.34 558.06
+Video–Audio Statistics
+Media Length 319.98 381.24 327.25 315.77 1039.18 289.09 222.42 259.69 352.63
+Media min. Length 6.15 10.03 5.57 5.57 10.04 5.57 5.57 10.15 5.57
+Media max. Length 6620.63 4427.76 3205.50 6620.63 6620.63 3851.93 1800.17 511.23 6620.63
+Media std. 527.50 681.63 482.65 603.49 1341.37 439.54 268.20 117.12 610.82
+Table 2: Statistics from the included data.
+Agentic Categories: counts per category.
+QA Lengths: mean word counts for questions, mean per-option length (MCQs), and open-ended answers (computed only when options are absent).
+Video–Audio Statistics: duration in seconds (mean, min, max, std).
+Quality Assurance
+To ensure the reliability of MAVERIX’s videos and annotations, each QA pair undergoes four checks by an expert annotator, as illustrated in Fig.
+3: (1) linguistic validity for clear and grammatical phrasing, (2) answerability of whether the question is resolvable via the video’s audiovisual content), (3) option integrity to ensure one correct answer with plausible distractors like semantically tangent or structurally identical options, and (4) modality interdependence, using cross-modal invalidation tests from Section Evaluation Protocol (e.g., disabling audio or video to detect shortcuts).
 For open-ended questions, reviewers also confirm that rephrased variants preserve meaning without overlapping with the ground-truth wording.
-To guard against textual or visual biases, we test each question with GPT-4o for text-only and video keyframes + subtitles inputs.
-Questions answered correctly without full audiovisual reasoning are revised.
-For example, if GPT-4o answers “Why did the mechanic sigh?” from subtitles alone, we refine the question.
-Reviewers document model failures and work with annotators to iteratively resolve ambiguous cases.
-This combined human-model process ensures MAVERIX’s questions demand proper multimodal understanding.
-2.2 MAVERIX Statistics
-MAVERIX comprises a diverse set of videos spanning 155 real-world scenarios across seven agentic task domains, such as social interactions, sentiment analysis, and egocentric understanding.
+Figure 3: The framework to construct annotation sets with hybrid annotator and MLLM-as-judge quality assurance.
+Dual‑Format Evaluation
+[…]
+Open‑Ended Generation.
+Every clip is paired with at least one semantically unique free-form prompt, ranging from causal explanation to future prediction, requiring natural‑language output.
+For each unique question, we paraphrase the question and prompt the model again to test its robustness against paraphrasing.
+The open-ended responses are graded by GPT‑4o on a five‑factor rubric covering factual correctness, attention to detail, contextual grounding, temporal coherence, and paraphrase robustness.
+Together, the MCQ and generation tracks marry scalability, through high-throughput accuracy metrics, with depth, by exposing weaknesses in explanation quality.
+MAVERIX Statistics
+MAVERIX comprises a diverse set of videos spanning 155 real-world scenarios across evaluation dimensions including agentic categories, topic domains, sub-topic domains, audio categories, duration, and difficulty.
 These are complemented by hierarchies over topics (e.g., travel, technology), video categories (e.g., documentaries, vlogs), and multimodal abilities (e.g., temporal reasoning, emotion recognition), supporting broad and balanced evaluation.
-Each question is also tagged with a difficulty level reflecting its multimodal complexity as judged by human annotators.
-A detailed breakdown of these splits is provided in Fig.
+Each question is also tagged with a difficulty level judged by human annotators.
+A detailed breakdown of these splits is provided in Tab.
+2.
+The dataset contains 105.8 hours of video footage, with durations ranging from 10 seconds to 63 minutes.
+Videos are distributed across three duration categories: 16.8% short clips (<1 minute) for rapid context-switching understanding, 75.7% medium-length videos (1-10 minutes) for sustained understanding, and 7.4% long-form content (10-65 minutes) for testing temporal coherence.
+The average video length is 352.63 seconds, with an std.
+of 610.82, suggesting that the included medias have a diverse distribution in lengths.
+Constructing a audiovisual benchmark with diverse media lengths reflects real-world use cases for MLLMs and sufficiently challenges their ability to generalize.
+Each video is paired with 3 to 4 questions on average with 2,556 in total, including 852 eight-option multiple-choice questions (MCQs) and 1704 open-ended prompts, with examples shown in Fig.
+2.
+Questions span over the evaluation dimensions to ensure a thorough evaluation.
+To mitigate positional bias, answer labels are uniformly redistributed across options.
+Experiments
+Evaluation Protocol
+MAVERIX adopts a dual evaluation framework to assess multimodal LLMs (MLLMs) through eight-option multiple-choice questions (MCQs) and open-ended response generation.
+Evaluation is conducted under two settings: localized, where models access only the timestamped video segment relevant to each question, and global (MAVERIX-Long), where the full-length video is provided.
+The localized setting limits the context to the specific temporal window required for understanding, whereas providing the full videos demands the models to localize the required information from the haystack of frames.
+For MCQs, we report both split-specific and overall accuracy, with answer choices uniformly distributed across positions (A-H) to reduce positional bias.
+Open-ended responses are evaluated using a GPT-4o scoring pipeline, adapted from Video ChatGPT (Maaz et al.
+2024), which assesses the output in five dimensions on a scale of 0-5.
+The results are aggregated across modalities (Tab.
+3), with separate analyses for easy, medium, and hard videos to diagnose comprehension limitations in Tab.
 4.
-The dataset contains 105.8 hours of video footage, with durations ranging from 10 seconds to 63 minutes.
-Videos are distributed across three duration categories: 16.8% short clips (<1 minute) for rapid context-switching tasks, 75.7% medium-length videos (1–10 minutes) for sustained narrative reasoning, and 7.4% long-form content (10–65 minutes) for testing temporal coherence.
-The average video length is 5.7 minutes, reflecting real-world use cases like troubleshooting tutorials or sports highlights.
-Each video is paired with 3–4 questions on average (2,556 total), including eight-option multiple-choice questions (MCQs) and open-ended prompts.
-Questions are evenly rebalanced across tasks to ensure a balanced evaluation.
-To mitigate positional bias, answer labels are uniformly redistributed across options (e.g., correct answers appear equally in positions A–H).
-Fig.
-4 details the dataset’s domain composition, and Fig.
-5 shows the duration distribution and question-type breakdown.
-3 Experiments
-3.1 Evaluation Protocol
-MAVERIX adopts a dual evaluation framework to assess multimodal LLMs (MLLMs) through eight-option multiple-choice questions (MCQs) and open-ended generation tasks.
-Evaluation is conducted under two settings: localized, where models access only the timestamped video segment relevant to each question, and global, where the full-length video is provided.
-The localized setting mitigates information leakage by limiting the context to the specific temporal window required for reasoning.
-For MCQs, we report both task-specific and overall accuracy, with answer choices uniformly distributed across positions (A–H) to reduce positional bias.
-Open-ended responses are evaluated using a GPT-4o-based scoring pipeline, adapted from Video-ChatGPT [35], which assesses outputs on five dimensions using a 0–5 scale.
-Scoring is calibrated on 200 human-annotated examples to ensure consistency, with penalties applied for failures in audiovisual reasoning, such as overlooking modality contradictions.
-Results are aggregated across tasks (Tab.
-1), with separate analyses for easy, medium, and hard videos to diagnose reasoning limitations.
-This protocol ensures reproducibility while addressing modality interdependence, which is the core challenge MAVERIX is designed to evaluate.
-3.2 Baselines
-We evaluate MAVERIX on a diverse suite of 12 MLLMs, encompassing both commercial and open-source architectures, to assess their ability to reason over intertwined audiovisual modalities.
-Commercial models include Gemini 1.5 Pro [49], GPT-4o [39], o1 [38], Grok2-Vision [52], Claude Sonnet 3.5 [5], Nova-Lite [23], and Nova-Pro [23], while open source representatives feature InternVL2 [14], Qwen2.5-VL [6], LLaVA-OneVision [27], DeepSeek-VL2-Tiny [51], and DeepSeek-VL2-Small [51].
-Gemini 1.5 Pro is uniquely equipped with native audiovisual processing, enabling direct ingestion of raw video-audio streams.
-For these models, we maximize temporal resolution by sampling frames at their maximum supported rates.
+We also report the model’s open-ended response qualities and token counts in Tab.
+6.
+Our proposed evaluation protocol ensures reproducibility while addressing modality interdependence and human performance baselines.
+Baselines
+We evaluate MAVERIX on a diverse suite of 17 MLLMs, encompassing both proprietary and open-source models, to assess their ability to reason over intertwined audiovisual modalities.
+Proprietary models include Gemini 2.0-FL (Team and Petko Georgiev 2025), Gemini 2.5-FL (Team and Petko Georgiev 2025), GPT-4o (OpenAI and Aaron Hurst 2024), Grok4 (xAI 2025), Claude Sonnet 3.5 (Anthropic 2024), Nova-Lite (Intelligence 2024), and Nova-Pro (Intelligence 2024).
+While open source representatives feature Ola (Liu et al.
+2025b), EgoGPT (Yang et al.
+2025), VITA 1.5 (Fu et al.
+2025), Qwen 2.5 Omni (Xu et al.
+2025), InternVL2 (Chen et al.
+2024c), Qwen2.5-VL (Bai et al.
+2025), LLaVA-OneVision (Li et al.
+2024a), DeepSeek-VL2-Small (Wu et al.
+2024).
+Among them, Ola, EgoGPT, VITA 1.5, Qwen 2.5 Omni and Gemini are equipped with native audiovisual processing, enabling direct ingestion of raw video-audio streams.
+For the tested models, we maximize temporal resolution by sampling frames at their maximum supported rates.
 However, most architectures, including GPT-4o and LLaVA-OneVision, require transcribed subtitles as text proxies for audio.
-To standardize inputs, we preprocess all videos using Whisper-base-v3 to extract time-synced subtitles, with the timestamps provided to the evaluated models.
+To standardize inputs, we preprocess all videos using Whisper-v3(Radford et al.
+2022) to extract time-synced subtitles, with the timestamps provided to the evaluated models.
+[…]
+Model
+Audio
+Support
+Size Arch Recipe Unimodal Analysis Multimodal Performance
+[…]
+Human $44.3$ $41.7$ 81.4 81.4 86.4 92.8 92.8 +11.4+11.4
+EgoGPT-7B 7B Dual-Tower SFT $29.9$ $43.0$ $45.2$ $45.2$ $55.0$ $45.2$ $55.0$ +9.8+9.8
+Ola-7B 7B Tri-Tower SFT 49.4 $43.9$ $37.6$ $49.5$ $49.6$ $53.1$ $53.1$ +3.6+3.6
+VITA 1.5 7B Dual-Tower SFT $32.4$ $43.4$ $20.2$ $43.5$ $22.3$ $18.5$ $22.3$ 21.2-21.2
+Qwen 2.5 Omni 7B Dual-Tower SFT+RL $46.5$ $41.4$ $35.4$ $46.5$ $57.9$ $49.5$ $57.9$ +11.4+11.4
+Qwen-2-VL 7B ViT-LLM SFT $43.0$ $48.0$ $48.0$ $57.5$ $57.5$ +9.5+9.5
+Qwen-2.5-VL 7B ViT-LLM SFT $40.3$ $46.9$ $46.9$ $55.3$ $55.3$ +8.4+8.4
+InternVL2 8B ViT-LLM SFT $24.1$ $26.3$ $26.3$ $33.1$ $33.1$ +6.8+6.8
+LLaVA-OneVision 7B SigLIP-LLM SFT $44.5$ $46.8$ $46.8$ $55.6$ $55.6$ +8.8+8.8
+DeepSeekVL2-small 2.8B Hybrid Enc.-MoE SFT $34.3$ $33.2$ $34.3$ $42.4$ $42.4$ +8.1+8.1
+Gemini 2.0-FL $43.8$ $38.0$ $42.1$ $43.8$ $41.1$ $50.2$ $50.2$ +6.4+6.4
+Gemini 2.5-FL $44.8$ $47.7$ $48.8$ $48.8$ $56.7$ $54.7$ $56.7$ +7.9+7.9
+Claude Sonnet 3.5 $55.0$ $42.0$ $55.0$ 64.1 64.1 +9.1+9.1
+GPT-4o 55.3 $54.3$ 55.3 $64.0$ $64.0$ +8.7+8.7
+Grok 4 $41.8$ $54.5$ $54.5$ $59.4$ $59.4$ +4.9+4.9
+GPT-4o-mini $45.4$ $35.5$ $45.4$ $50.0$ $50.0$ +4.6+4.6
+NOVA-Lite $40.4$ $39.7$ $40.4$ $51.0$ $51.0$ +10.6+10.6
+NOVA-Pro $46.6$ $45.4$ $46.6$ $55.8$ $55.8$ +9.2+9.2
 All models receive inputs in the unified format [video frames, subtitles, question], with frames uniformly sampled at their maximum supported context window.
-For audio-incapable models, subtitles replace raw audio tracks, while Gemini 1.5 Pro processes synchronized audio-video pairs.
+For audio-incapable models, subtitles replace raw audio tracks, while the audio-supported models additionally process synchronized audio-video pairs.
 We employ a standardized prompt template across models, ensuring fairness by eliminating instructional biases.
-This setup isolates modality interdependence as the critical challenge: models must synthesize visual cues like the startled expression of a character with an auditory context like an off-screen crash to match human-like reasoning.
-3.3 Results
-Models Params Easy (%) Medium (%) Hard (%) Overall (%)
-w/o sub w/ sub w/o sub w/ sub w/o sub w/ sub w/o sub w/ sub
-Random Guess - 12.4 12.7 12.3 12.5
-Human (Video) - 60.2 48.8 59.4 55.3
-Human (Audio) - 34.9 27.4 40.8 31.1
-Human (Video+Audio) - 83.6 76.2 84.4 80.3
-Open-source MLLMs (Truncated Videos)
-InternVL2 [14] 2B 31.1 35.2 (+4.1) 27.7 35.1 (+7.4) 17.8 24.3 (+6.5) 27.5 33.4 (+5.9)
-Qwen2.5-VL [6] 7B 30.8 34.9 (+4.1) 25.9 31.9 (+6.0) 19.3 20.7 (+1.4) 26.8 31.3 (+4.5)
-LLaVA-OneVision [27] 7B 26.6 28.9 (+2.3) 26.5 28.2 (+1.7) 29.9 29.9 (+0.0) 27.1 28.8 (+1.7)
-DeepSeek-VL2-Tiny [51] 3.37B 16.8 15.1 (-1.7) 14.5 20.6 (+6.1) 20.5 17.3 (-3.2) 16.4 17.9 (+1.5)
-DeepSeek-VL2-Small [51] 16.1B 39.9 46.4 (+6.5) 38.8 46.4 (+7.6) 25.9 29.6 (+3.7) 37.2 43.8 (+6.6)
-Closed-source MLLMs (Truncated Videos)
-GPT-4o [39] - 58.9 69.8 (+10.9) 60.7 67.0 (+6.3) 41.5 49.6 (+8.1) 56.9 65.4 (+8.5)
-o1 [38] - 74.0 79.3 (+5.3) 70.7 76.0 (+5.3) 55.6 67.4 (+11.8) 69.6 75.9 (+6.3)
-Claude Sonnet 3.5 [5] - 27.4 68.9 (+41.5) 39.6 64.6 (+25.0) 27.4 44.4 (+17.0) 40.3 63.1 (+22.8)
-Grok2-Vision [52] - 58.3 65.1 (+6.8) 55.1 61.2 (+6.1) 45.9 47.4 (+1.5) 54.9 60.6 (+5.7)
-Nova-Lite [23] - 51.8 59.8 (+8.0) 49.3 55.9 (+6.6) 37.8 43.0 (+5.2) 48.5 55.4 (+6.9)
-Nova-Pro [23] - 53.0 51.5 (-1.5) 51.7 47.5 (-4.2) 35.6 34.1 (-1.5) 49.6 46.9 (-2.7)
-Gemini 1.5 Pro [49] - 64.5 69.5 (+5.0) 65.4 69.4 (+4.0) 55.6 49.6 (-6.0) 63.5 66.3 (+2.8)
-Open-source MLLMs (Full-Length Videos)
-InternVL2 [14] 2B 23.0 26.3 (+3.3) 18.9 24.4 (+5.5) 14.2 11.0 (-3.2) 19.7 23.0 (+3.3)
-Qwen2.5-VL [6] 7B 47.4 55.9 (+8.5) 45.3 51.7 (+6.4) 37.9 47.2 (+9.3) 44.8 52.6 (+7.8)
-LLaVA-OneVision [27] 7B 34.5 40.1 (+5.6) 34.3 39.5 (+5.2) 30.7 35.4 (+4.7) 33.8 39.1 (+5.3)
-DeepSeek-VL2-Tiny [51] 3.37B 17.1 26.0 (+8.9) 15.4 21.5 (+6.1) 21.3 19.7 (-1.6) 17.0 23.0 (+6.0)
-DeepSeek-VL2-Small [51] 16.1B 28.3 44.1 (+15.8) 29.9 41.6 (+11.7) 23.6 26.0 (+2.4) 28.3 40.0 (+11.7)
-Closed-source MLLMs (Full-Length Videos)
-GPT-4o [39] - 57.6 64.8 (+7.2) 58.4 62.2 (+3.8) 43.3 46.5 (+3.2) 55.6 60.6 (+5.0)
-o1 [38] - 72.4 73.0 (+0.6) 68.3 72.4 (+4.1) 52.0 65.4 (+13.4) 67.2 71.5 (+4.3)
-Claude Sonnet 3.5 [5] - 7.9 17.4 (+9.5) 8.1 14.8 (+6.7) 7.9 17.3 (+9.4) 8.0 16.3 (+8.3)
-Grok2-Vision [52] - 55.6 59.2 (+3.6) 55.8 55.2 (-0.6) 45.7 53.5 (+7.8) 54.1 56.5 (+2.4)
-Nova-Lite [23] - 47.4 59.5 (+12.1) 45.3 55.9 (+10.6) 37.8 43.0 (+5.2) 44.9 55.3 (+10.4)
-Nova-Pro [23] - 45.2 58.2 (+13.0) 48.4 58.1 (+9.7) 28.9 41.7 (+12.8) 44.1 55.5 (+11.4)
-Gemini 1.5 Pro [49] - 65.1 69.4 (+4.3) 64.0 68.0 (+4.0) 54.3 55.1 (+0.8) 62.8 66.5 (+3.7)
-Table 1: Performance comparison of different MLLMs across various settings with and without subtitles.
-Video only Audio only Text only Video + Sub Video + Audio
-Truncated Full Truncated Full Truncated Full Truncated Full Truncated Full
-Level Easy 64.5 65.1 56.8 57.6 45.6 51.6 69.5 69.4 76.0 74.0
-Medium 65.4 64.0 57.8 52.3 47.0 46.8 69.4 68.0 71.5 70.3
-Hard 55.6 54.3 46.7 49.6 34.8 41.7 49.6 55.1 63.0 62.2
-Agentic Tasks Social 55.6 58.3 47.5 49.0 41.4 54.2 62.6 59.4 61.6 67.7
-Sentiment 69.1 70.4 53.1 51.9 34.6 42.0 66.7 64.2 76.5 74.1
-Egocentric Agent 64.4 61.7 49.2 53.2 33.9 31.9 57.6 55.3 78.0 68.1
-Information Querying 66.8 65.9 58.9 55.6 48.3 50.0 69.4 68.3 75.1 72.2
-Sports 49.6 50.5 54.6 47.6 45.4 45.6 64.7 66.0 63.9 62.1
-Gaming 63.1 61.2 53.4 49.4 41.7 43.5 58.3 64.7 63.1 62.4
-Shopping 72.2 67.6 62.7 65.8 51.6 54.1 74.6 76.6 82.5 81.1
-Audio Type Natural Sound 57.4 64.6 47.1 50.8 29.4 32.3 54.4 61.5 64.7 61.5
-Speech 63.5 61.4 58.7 55.2 51.6 55.6 69.3 70.3 76.2 73.1
-Music 69.2 72.0 53.8 56.0 7.7 24.0 69.2 56.0 76.9 76.0
-Artificial Sound 60.0 74.0 36.0 40.0 20.0 26.0 56.0 58.0 56.0 68.0
-Mixed Sounds 65.3 61.5 56.5 55.5 44.7 45.4 66.0 63.8 69.1 67.9
-Topics Humanities and Society 70.7 71.3 63.9 65.9 48.9 51.9 77.4 79.1 82.7 80.6
-Geography and Travel 68.8 68.8 56.2 56.2 62.5 68.8 50.0 68.8 75.0 75.0
-Technology and Gaming 64.9 64.0 52.7 49.1 42.0 41.2 61.1 62.3 65.6 64.9
-Science and Knowledge 57.1 63.0 39.3 51.9 39.3 37.0 53.6 63.0 67.9 70.4
-Movies, TV and Animations 61.2 61.2 35.8 37.3 29.9 44.8 55.2 59.7 59.7 65.7
-Arts and Performance 58.6 60.7 48.3 42.9 34.5 39.3 58.6 57.1 62.1 71.4
-Pets and Animals 64.4 66.7 55.6 48.9 40.0 44.4 64.4 60.0 71.1 73.3
-Business and Commerce 73.3 68.6 68.6 71.4 55.8 57.1 79.1 77.1 86.0 87.1
-Life and Practical Skills 72.1 64.7 69.2 60.0 54.8 58.8 77.9 64.7 84.6 77.9
-Sports and Adventure 53.2 56.9 52.4 48.6 43.7 45.9 65.1 69.7 63.5 63.3
-Social Trends and Reactions 51.9 50.0 40.7 46.2 38.9 48.1 57.4 57.7 59.3 55.8
-Vehicles and Transportation 55.6 55.6 55.6 44.4 22.2 11.1 55.6 44.4 66.7 55.6
-Low Quality and Extended Content 41.7 41.7 54.2 54.2 29.2 37.5 37.5 50.0 66.7 62.5
-Multimodal Abilities Visual Reasoning 63.2 65.7 54.8 57.4 43.4 46.6 64.5 64.2 72.8 71.1
-Emotional Inference 67.4 68.3 55.8 51.2 37.2 34.1 62.8 58.5 72.1 73.2
-Situational Reasoning 70.0 77.8 55.0 50.0 50.0 44.4 70.0 66.7 80.0 77.8
-Causal Reasoning 65.7 60.3 47.8 43.1 37.3 48.3 67.2 65.5 70.1 60.3
-Factual Recall 64.3 61.0 57.9 55.1 46.2 48.6 68.2 69.0 72.5 71.7
-Spatial Reasoning 37.9 45.8 51.7 33.3 41.4 41.7 58.6 45.8 55.2 50.0
-Relationship 62.5 62.5 43.8 62.5 50.0 75.0 62.5 75.0 62.5 75.0
-Temporal Distance 66.7 81.8 58.3 54.5 58.3 72.7 58.3 81.8 83.3 81.8
-Overall 63.5 62.8 55.6 53.9 44.5 47.9 66.3 66.5 71.9 70.5
-Table 2: Performance comparison of Gemini 1.5 Pro across different modalities, difficulties, and categories.
-We select a wide range of open-sourced and proprietary MLLMs on the MAVERIX benchmark.
-We report the quantitative results grouped by difficulties in Tab.
-1 for their performances with and without subtitles and by different video lengths.
-The detailed quantitative GPT consistency scores can by found in Tab.
-D.2, and visualized in Fig.
-6.
-We then showcase the fine-grained quantitative evaluations of Gemini 1.5 Pro in different multimodal input settings and grouped by the splits in Tab.
-2.
-Refer to Section D for additional experimental results.
-Below, we provide our key observations.
-Figure 6: Consistency of model responses on open-ended questions, measuring how reliably models provide coherent and relevant answers across different inputs.
+This setup isolates modality interdependence as the critical challenge: models must synthesize potentially asynchronous audiovisual cues, such as the startled expression of a character with an auditory context like an off-screen crash to match human-like understanding.
+Table 4: Difficulty-wise MCQ accuracy (%) for audio-enabled models. AV cells show $\Delta$ vs A.
+Model Easy Medium Hard
+A V AV A V AV A V AV
 […]
@@
 […]
-Are Audio and Text Equally Informative?
-Refer to Tab.
-2 for detailed results.
-We evaluated Gemini 1.5 Pro, which was selected for its support for audio, text, and video modalities.
-Across all categories, Gemini 1.5 Pro consistently achieves higher scores with audio-only input compared to text-only input, suggesting that the model may extract more informative or discriminative cues from audio signals than from textual descriptions in this setting.
-Furthermore, when comparing multimodal configurations, video+audio yields the highest overall performance, surpassing video+subtitle.
-Although video + audio generally provide the most significant benefit, certain categories, including sentiment understanding, situational reasoning, causal reasoning, and egocentric agent tasks, experience particularly large performance gains under this configuration.
-These tasks likely benefit from the richer contextual and temporal cues present in audio, such as tone, ambient sounds, and speaker intent, which are difficult to fully capture through visual or textual modalities alone.
-Misaligned Category Trends Between MLLMs and Humans: The category-level analysis in Fig.
-1 reveals notable differences in how MLLMs and humans perform across tasks.
-The quantitative comparisons are detailed in Tab.
-D.6.
-Humans tend to excel in dynamic contextual categories such as egocentric reasoning, sports analysis, and gaming strategy, which require interpreting spatial cues, tracking actions over time, and understanding complex interactions.
-However, their performance is less consistent in domains like shopping and social interaction.
-Shopping tasks may prioritize factual or procedural details that are less engaging or familiar to humans.
-For social interaction tasks, inherent subjectivity, such as interpreting cultural contexts, intentions, or unspoken tensions, can lead to diverse and sometimes conflicting human judgments, reducing overall precision.
-MLLMs show a different trend.
-They perform comparatively well on shopping-related tasks, which often involve concrete retrieval of explicit audiovisual details like product descriptions, prices, or transactional cues.
-These structured object-centric queries align well with the strengths of large models trained on web-scale corpora rich in product and factual information.
-In contrast, MLLMs consistently struggle in categories that require nuanced temporal reasoning, social understanding, and dynamic context integration, such as predicting team strategies in games or recognizing subtle emotional cues in conversation.
-These trends suggest that current MLLMs are more adept at tasks grounded in static, factual associations but remain limited in handling complex, situational reasoning that requires modeling interactions, intent, and evolving context.
-Closing this gap will likely require advances in temporal modeling, sentiment cue understanding, and the ability to generalize beyond explicit audiovisual patterns.
-4 Related Work
-MLLM Benchmarks: The development of MLLM benchmarks has progressed from static image-based question answering (QA) to dynamic video understanding.
-Early benchmarks for evaluating vision-language models (VLMs) focused primarily on images, emphasizing visual grounding and object recognition.
-Earlier datasets  [11, 3, 59, 41] provided human-annotated captions and QA pairs to assess these core capabilities.
-Building on this foundation, subsequent benchmarks [42, 34, 34, 44] introduced more domain-specific and knowledge-intensive tasks.
-A-OKVQA [43] challenges abilities to read and interpret texts within images.
-More recently, MMMU [54] and MMMU-Pro [55] have expanded these efforts with diverse tasks designed to probe a model’s ability to simultaneously ‘see” and read” from prompted images.
-Despite their advances, image-based benchmarks lack the temporal and acoustic context necessary for richer multimodal understanding.
-This limitation motivated the shift toward video-based evaluation.
-Fine-grained temporal tasks assessing a model’s ability to perceive motion and temporal sequences have been included in video benchmarks [29, 16, 40, 25, 22, 45, 35, 30].
-However, these datasets and recent extensions  [37, 12, 21, 36, 46] tend to feature short clips with constrained domains and limited semantic diversity.
-They primarily probe local temporal, visual semantics, and factual information, offering limited insight into how models handle the complexities of daily activities, situational awareness, or social reasoning.
-Most recent works [9, 29, 40, 25, 30, 10] take a step towards everyday contexts by incorporating human actions, but their scopes are restricted to human-centric activities and do not handle visual-audio cues.
-Video-MME [17] is a comprehensive video benchmark with short, medium, and long videos.
-However, it only employed the MCQ format, overlooking the quality and richness of free-form expressions generated by MLLMs from open-ended questions.
-MAVERIX extends prior benchmarks by explicitly elevating audio as a primary modality, with its questions requiring complex cross-modal reasoning, such as resolving conflicts between dialogue and visual cues.
-Unlike existing datasets, it systematically evaluates a model’s ability to integrate complete multimodal signals through both descriptive open-ended answers as well as multiple-choice questions, comparisons shown in Tab.
-3.
-Video Understanding Models: Models based on contrastive pre-training over image-text pairs have demonstrated significant efficacy in learning transferable representations across various visual tasks.
-These methods have been adapted to specialized domains, including multi-disciplinary scientific understanding and multimodal reasoning.
-Building upon advancements in pre-training, subsequent work [50, 6, 48, 56, 57, 32] has sought to enhance the multimodal reasoning capabilities of complex spatio-temporal video data and improve their alignments with human preferences through post-training reinforcement learning from human feedback (RLHF).
-Other efforts [51, 31, 15, 47, 7] integrating a mixture of experts (MoE) architectures [33], which conditionally activates subsets of specialized “expert” networks during processing, have shown improved scalability and performance without a proportional increase in computational costs.
-However, whether these developments enable models to reason in ways comparable to or surpass humans remains an open question [1, 4, 8, 24], as the assessment of human-like multimodal reasoning, particularly in real-world visual-audio contexts, remains underexplored.
 […]
-Benchmarks #Videos Len. #Questions Subtitle Audio Open Multi-Split
-MSRVTT-QA 2,990 15.2 72,821
-MSVD-QA 504 9.2 50,505
-TGIF-QA 9,575 3.6 165,165
-ActivityNet-QA 5,800 14.9 58,000
-TVQA 2,179 15.2 152,545
-How2QA 1,517 11.5 71,812
-STAR 914 11.9 59,871
-NExT-QA 4,269 12.1 52,631
 […]
-Video-Bench 5,917 56.0 17,000
-EgoSchema 6,365 18.0 36,054
-AutoEval-Video 327 32.1 450
-TempCompass 410 11.4 1,540
-Video-MME 900 1071.9 2,700
-HourVideo 500 2742 12,976
 […]
-MAVERIX 700 344.9 2,556
-Table 3: Comparison of video QA benchmarks with various attributes.
-5 Conclusion
-Agentic tasks such as assisting collaborative work and navigating dynamic environments require strong audiovisual reasoning capabilities, and yet these abilities have not been fully assessed in the recent MLLMs.
-To address this, we introduce MAVERIX, a benchmark designed to evaluate models in complex real-world audiovisual scenarios.
-It contains 700 videos and 2,533 human carefully crafted questions providing not only a challenging benchmark but also a comprehensive evaluation suite.
-We conduct a thorough benchmarking of mainstream closed-source and open-source MLLMs on the proposed benchmark.
-Our evaluations reveal that cross-modal video-audio understanding is essential for solving real-world perception problems.
-Additionally, they show that even state-of-the-art MLLMs still exhibit a performance gap compared to humans, particularly in socially grounded or dynamic tasks.
-We hope MAVERIX will serve as a foundation for advancing research in multimodal reasoning and guiding future work towards more context-aware and socially intelligent systems.
-Supplementary Material
-MAVERIX Supplementary Material
-  • Section A: MAVERIX Release v1.0
-  • Section B: Rationale and Liminations
-  • Section C: Qualitative Comparisons by Modalities
-  • Section D: Additional Experiments
-    • Section D.1: Standardized Prompt Design
-    • Section D.2: Performance by Difficulties
-    • Section D.4: Quantitative GPT Evaluations
-    • Section D.5: Performance by Categories
-    • Section D.6: Common Error Modes
-  • Section E: Details on the Human Survey
-  • Section F: Additional Data Distribution Analysis
-Appendix A MAVERIX Release v1.0
-We are releasing MAVERIX v1.0, our proposed benchmark dataset for multimodal video-language understanding.
+[…]
+Model analysis with Maverix
+This section discusses model behaviour along five dimensions: multimodal gains across modal deisgns, training‑recipe variation, agentic ability relative to humans, temporal‑horizon sensitivity, and perceived question difficulty, using the benchmark set.
+Multimodal Gains Across Model Capacity and Architecture.
+Tab.
+3 summarizes performance across five architectural families and multiple model sizes.
+Best unimodal accuracies span $\sim$26–55%, and most models show gains when additional modalities are provided.
+Many systems follow an audiovisual encoder++MLP++LLM design; among these, Ola, EgoGPT, and Qwen 2.5 Omni, which incorporate Whisper v3 for audio, generally do not regress and often improve relative to their strongest unimodal scores, whereas VITA 1.5 shows a regression with audio–visual input.
+The lightweight Gemini 2.0-FL and 2.5-FL variants also improve with multimodal inputs.
+For video-text models, adding a modality yields consistent gains; Qwen2-VL, Nova-Lite, Nova-Pro, and Claude 3.5 Sonnet improve by roughly $\sim$10%.
+Despite these gains, a sizable gap to human performance remains, suggesting that current models under-utilize cross-modal cues.
+This is evident when contrasting V++S and V++A: several models (e.g., Gemini 2.5-FL, EgoGPT, VITA 1.5, and Qwen 2.5 Omni) score lower with V++A than with V++S, indicating missed auditory details or limitations in audio-video fusion.
+Table 5: Accuracy (%) on AV inputs across taxonomy and audio categories (single-column).
+Taxonomy (AV)
+Models Social Ego-centric Sports Gaming
+Human 92.7 95.2 82.3 74.5
+EgoGPT 34.3 45.8 42.9 40.8
+Ola 52.5 57.6 52.9 46.6
+VITA 1.5 21.2 6.8 16.8 12.6
+Qwen 2.5 Omni 47.5 42.4 49.6 41.7
+Gemini 2.0-FL 46.5 32.3 48.7 47.6
+Gemini 2.5-FL 47.1 39.0 54.3 61.9
+Audio Category (AV)
+Models Natural Speech Music Artificial Mixed
+Human 88.4 94.1 90.3 89.6 92.8
+EgoGPT 41.2 44.2 46.2 48.0 47.3
+Ola 48.5 52.0 38.5 58.0 56.5
+VITA 1.5 29.4 17.0 7.7 26.0 17.6
+Qwen 2.5 Omni 52.9 47.8 53.8 48.0 51.5
+Gemini 2.0-FL 47.1 51.1 38.5 54.0 49.6
+Gemini 2.5-FL 51.5 56.7 50.0 52.0 53.1
+Training-Recipe Variants: SFT, RL, and Data Composition.
+We assess open-source recipes built on identical backbones to isolate curriculum effects.
+Omni-modal models such as Ola and Qwen 2.5 Omni use an image-text warmup followed by separate alignment for the audio modality, and they typically improve on audiovisual evaluations; reinforcement learning appears to further increase the multimodal gains.
+VITA 1.5, on the other hand, emphasizes alignment to the video modality during training, which may bias attention toward the visual stream and results in regression when subtitles and audio are added.
+Video-text models follow a more streamlined path: initial pretraining on image-caption pairs to align images with text, then instruction tuning or long chain-of-thought data for fine-tuning.
+Aside from minor architectural differences, they vary primarily in data curation and sources.
+Qwen 2.5 VL uniquely includes chain-of-thought data during fine-tuning to encourage explicit reasoning and stronger multimodal synthesis.
+While its overall performance is strong, the relative gain from multimodal inputs appears similar to its counterpart without chain-of-thought fine-tuning, which may point to reward hacking during SFT and warrants further study.
+Table 6: Open-ended response correctness scores (out of 5) as judged by GPT, reported per modality.
+For models without native audio support, scores for Subtitle (S) and Subtitle+Video (SV) are shown instead of Audio (A) and Audio+Video (AV).
+Response token length statistics are analyzed separately.
+Model
+Audio
+Support
+Judged Score Token Length
+A/S V AV/SV Avg Max Std
+[…]
+[…]
+[…]
+[…]
+Qwen-2-Omni 1.0 1.2 1.2 $0\downarrow$ 47.0 178 26.5
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Agentic Ability in Comparison with Humans.
+Tab.
+5 indicates that humans perform best on social and egocentric questions, with slightly lower accuracy on gaming and sports that demand domain knowledge.
+Model behavior is less uniform.
+Gemini, Qwen 2.5 Omni, and VITA 1.5 tend to be weaker on egocentric videos and comparatively stronger on domain-specific categories, while the remaining models show different per-category strengths.
+Across audio categories, human accuracy is largely stable.
+Models trained on broader multimodal corpora such as Qwen 2.5 Omni and the Gemini family exhibit smaller fluctuations across audio types, yet many systems under-perform when music dominates.
+EgoGPT shows relatively strong auditory understanding, likely reflecting its use of a pretrained Whisper encoder.
+In contrast, Ola and VITA 1.5 that trained on smaller datasets, display larger variance across categories, with notable drops on the music category.
+Taken together, these patterns suggest that coverage of first-person content and diverse audio during training may be as important as scale for robust agentic ability across domains.
+Temporal Horizons: Long- versus Short-Clip Performance.
+As shown in Fig.
+4, we evaluate models on short, pre-localized audiovisual clips in MAVERIX and on their full-length counterparts in MAVERIX-Long.
+Across models, localized clips yield higher accuracy.
+Among the agentic categories, the questions that depend on immediate, synchronous audiovisual cues, such as those from factual recall and near-term causal inference, show the smallest degradation.
+When the relevant segment is pre-localized, models can more reliably extract the necessary information.
+By contrast, social relationship, emotion, and situational understanding often rely on fine-grained and sometimes asynchronous cues distributed over time.
+Performance drops more on long videos, reflecting challenges in localizing these signals and integrating them over extended context.
+Overall, a gap to human performance remains, especially for longer videos and for recognizing subtle contextual cues.
+These trends suggest that current MLLMs are stronger at retrieving salient, object or event level signals than at integrating evolving context and social nuance over time.
+Figure 4: Impact of Video Length on Agentic Category Performances.
+(a) Accuracy change (%) from short to full-length videos across models and questions.
+(b) Accuracy gap (%) to human performance for short and full inputs.
+Model and Human Perception of Difficulty.
+We analyze performance by difficulty and observe that multimodal inputs often help most on easy items, with smaller gains on hard ones, though trends vary by model.
+For the audio-enabled models where split statistics are available, Gemini 2.0-FL improves by ++12.4% on easy items and ++3.3% on hard items, and Gemini 2.5-FL improves by ++11.5% on easy items and ++5.3% on hard items.
+Qwen 2.5 Omni shows a different pattern with substantial benefit on hard items as well.
+These mixed results suggest that current systems leverage straightforward cross-modal cues more reliably than they integrate sparse or subtle signals in harder cases.
+Although MAVERIX is designed to elicit cross-modal reasoning, some models still achieve moderate scores with a single modality, likely because many real videos contain aligned audio and visual streams that allow plausible inferences from partial evidence.
+GPT-4o-mini is one such example of respectable unimodal performance.
+Humans also benefit from aligned cues and strong priors, yet the jump from 81.4% with video-only to 92.8% with audiovisual highlights the value of genuine cross-modal understanding and sets a clear target for future modeling.
+We hope continued progress in cross-modal alignment will narrow this gap and eventually surpass the current human baseline.
+Related Work
+MLLM Benchmarks.
+Early vision-language benchmarks centered on images for grounding and recognition, using captions and QA pairs (Chen et al.
+2015; Agrawal et al.
+2019; Zhu et al.
+2016; Plummer et al.
+2015), followed by domain-specific and knowledge-intensive settings (Saikh et al.
+2022; Lu et al.
+2023; Singh et al.
+2019).
+A-OKVQA targets external-knowledge reasoning beyond visible content (Schwenk et al.
+2022).
+More recent efforts, including MMMU and MMMU-Pro, broaden question diversity and reading-from-image skills (Yue et al.
+2024a, b).
+Image-only evaluation lacks temporal and acoustic context, motivating video benchmarks that probe motion, sequence, and temporal localization (Li et al.
+2024b; Patraucean et al.
+2023; Kesen et al.
+2023; Huang et al.
+2020; Song et al.
+2024a; Maaz et al.
+2024; Fang et al.
+2025; Li et al.
+2024c; Ning et al.
+2023; Chen et al.
+2024a; He et al.
+2024; Mangalam, Akshulakov, and Malik 2023; Song et al.
+2024b).
+However, most emphasize short clips and constrained domains, rely on MCQ-only protocols, and provide limited coverage of everyday social or situational reasoning.
+Video-MME and AV-Odyssey scale video duration but remain MCQ-only, omitting open-ended assessment (Fu et al.
+2024; Gong et al.
+2024).
+MAVERIX elevates audio as one of the primary signals and stresses cross-modal integration as it evaluates both 8-way MCQs and open-ended responses to measure multimodal synthesis under realistic audiovisual conditions.
+Video Understanding Models.
+Contrastive pretraining on image-text data yields transferable representations and has been adapted to spatio-temporal reasoning; post-training with instruction tuning and RLHF further aligns models with human preferences (Sun et al.
+2023; Zhai et al.
+2024; Lin et al.
+2023; Wang et al.
+2024; Bai et al.
+2025; Zhang 2024).
+MoE-style routing improves scalability without linear cost growth (Wu et al.
+2024; Lin et al.
+2024; DeepSeek-AI and Aixin Liu 2025; Sun, Chen, and Yiqing Huang 2024; Cai et al.
+2024; Liu et al.
+2025a).
+Yet whether these advances enable human-comparable multimodal reasoning in real-world audiovisual settings remains open (tes 2024; Amirizaniani et al.
+2024; Campbell et al.
+2024; Kazemi et al.
+2024).
+Our evaluations on MAVERIX show a substantial gap to human accuracy on MCQs, often on the order of several tens of percentage points, and highlight persistent challenges in integrating temporal, social, and auditory cues in models with different architecture and training recipes.
+Conclusion
+Agentic scenarios such as assisting collaborative work and navigating dynamic environments require strong audiovisual reasoning, yet these abilities remain under-assessed in recent MLLMs.
+We introduce MAVERIX, a benchmark for complex, real-world audiovisual understanding, comprising 700 videos and 2,556 carefully crafted, human-authored questions.
+The suite evaluates models with both 8-way multiple-choice and open-ended responses.
+Our results indicate that multimodal inputs generally improve accuracy, but sizable gaps to human performance persist, especially for socially grounded or dynamic scenarios.
+Models benefit most when relevant segments are pre-localized and tend to struggle on longer videos that demand integrating subtle, asynchronous cues over time; robust audio integration also remains uneven across systems.
+We hope MAVERIX will guide progress toward stronger cross-modal alignment, better temporal reasoning, and more context-aware, socially intelligent models.
+Appendix A A. MAVERIX Release v1.0
+We are releasing MAVERIX v1.0, our proposed benchmark dataset for multimodal video-language understanding, built with support to a common benchmark platform LMMs-Eval (Li* et al.
+2024).
 The benchmark dataset is provided as a single JSON file for ease of use and straightforward integration with existing benchmarking pipelines.
@@
 Relevant timestamps are provided to allow precise video trimming.
-We provide all 700 video clips used in our benchmark, and users can download the corresponding videos from the Ego4D website after reviewing and accepting the Ego4D license agreement.
+We will provide all 1400 video clips used in our benchmark through HuggingFace (hug ), with both localized and full length videos, and users can download the corresponding videos from the Ego4D website after reviewing and accepting the Ego4D license agreement.
 We also provide sample videos with annotations from MAVERIX.
 All materials are available through the project page upon acceptance.
+The code snippets providing main input prompts, response aggregation, judge response generations are provided as part of the supplement materials, with unit test functions for testing data corruption included.
 Structure: MAVERIX v1.0 release is organized as follows:
@@
     • vid_only/: Contains all muted videos for evaluations without audio modality.
-    • vid/: Contains all videos for evaluations without audio modality.
+    • vid/: Contains all videos for evaluations with both video and audio modality.
+    • audio/: Contains all audio files for evaluations with audio modality.
     • subtitles/: Contains all subtitles for evaluations requiring subtitle access.
@@
   • src/
-    • video_utils.py: A script for dataloading, processing, and evaluation functionalities.
+    • maverix_benchmark.py: A script for dataloading, processing, and evaluation functionalities.
     • _default_template_yaml: A file to specify the default processing template for the evaluation. The data loading path, cache directory, and evaluation modes can be specified here.
-    • MAVERIXXX_dev: Individual setting files specifying the test settings for different modes. The currently supported modes are: audio_only, video_only, text_only, video_text, and video_text_audio.
-    • Remark: Except for the open-sourced models listed in the paper or available through LMMs-Eval[58, 26], all other experiments require access to proprietary models.
+    • maverix_<mode>.yaml: Individual setting files specifying the test settings for different modes. The currently supported modes are: audio_only, video_only, sub_only, video_sub, and video_audio.
+    • Remark: Except for the open-sourced models listed in the paper or available through LMMs-Eval(Zhang et al. 2024; Li* et al. 2024), all other experiments require access to proprietary models.
 Documentation:
@@
   • We will maintain MAVERIX, with all updates and new versions announced publicly.
-Appendix B Rationale and limitations
+Appendix B B. Rationale and limitations
 Broader Impact.
@@
 Despite these limitations, MAVERIX offers a step toward a more comprehensive and realistic evaluation of multimodal LLMs, with the goal of driving future work in building models capable of deeper, contextually grounded, and socially aware reasoning.
-Appendix C Qualitative Comparisons by Modalities
-We conduct a comprehensive analysis of the performance of Gemini 1.5 Pro, providing a breakdown across each custom-defined split and input modality to better understand its strengths and limitations throughout the benchmark.
-Specifically, we evaluate the model under different input configurations like Audio Only, Video Only, and Audio+++Video, as well as across diverse content-based splits covering topics, agentic task types, audio characteristics, and difficulty levels.
-The corresponding results are shown in Fig.
-C.1 for modality comparison, Fig.
-C.2 for topic-wise performance, Fig.
-C.3 for agentic task types, Fig.
-C.4 for audio type variation, and Fig.
-C.5 for difficulty levels.
-This detailed analysis highlights how well Gemini 1.5 Pro generalizes across different challenge dimensions and input conditions.
-Figure C.1: Evaluation of multimodal abilities by Categories.
-Figure C.2: Evaluation of multimodal abilities by Topics.
-Figure C.3: Evaluation of multimodal abilities by Agentic Tasks.
-Figure C.4: Evaluation of multimodal abilities by Audio Types.
-Figure C.5: Evaluation of multimodal abilities by the level of difficulties.
-Appendix D Additional experiments
+Appendix C C. Additional experiments
 This section presents extended experimental results to further analyze model performance across various dimensions of the benchmark.
 In addition to model evaluation, we provide human performance baselines segmented by difficulty levels, input modalities, and agentic task categories for direct comparison.
-D.1 Standardized Prompt Design
+C.1 Standardized Prompt Design
 To ensure fairness in evaluation, we adopt a standardized prompt format across all models.
@@
 For tasks requiring multiple modalities, the prompts incorporate the modalities in the inputs.
-In multiple-choice settings, the prompt is presented as follows: {mdframed} [images] + [audio] (if available) This video’s subtitles are listed below: OR No subtitles available.
+In multiple-choice settings, the prompt is presented as follows: [images] + [audio] (if available) This video’s subtitles are listed below: OR No subtitles available.
 [Subtitles] Select the best answer to the following multiple-choice question based on the video and the subtitles.
 Respond with only the letter (A, B, C, D, E, F, G, or H) of the correct option.
-Respond with only the letter (A, B, C, or D) of the correct option.
 [Question] The best answer is:
 For open-ended questions, the following prompt is used:
-{mdframed}
 [images] + [audio] (if available) This video’s subtitles are listed below: OR No subtitles available.
@@
 [Question] The best answer is:
-D.2 Model Performance by Difficulties
-We evaluate model performance across different difficulty levels to assess robustness and generalization from simpler to more complex scenarios.
-The results are presented in Figure D.1, which illustrates how the model’s accuracy declines as task complexity increases and showcases its relative performance compared to human participants.
-Figure D.1: Probed model performance across different difficulty levels.
-This analysis illustrates how the models’ accuracy changes with increasing task complexity and their relative positions against humans.
-D.3 Human Performance by Categories
-We report performance across distinct agentic task categories to showcase the strengths and weaknesses of the human study participants within the defined contexts.
-We note that these statistics may be biased due to the composition of the participant pool, and further details on the recruitment process can be found in Sec.
-E.
+C.2 GPT-4o as Judge Criteria
+For evaluating the open-ended answer qualities, we divide the evaluations in 4 different aspects: correctness, level of details, context understanding, and temporal consistency.
+We prompt the judge model as follows for correctness:
+[…]
+The template prompt for level of detail orientation:
+[…]
+The template prompt for level of context understanding:
+[…]
+[…]
+C.3 Human Performance by Categories
+We report performance across distinct agentic task categories in Tab.
+7 to showcase the strengths and weaknesses of the human study participants within the defined contexts.
+We note that these statistics may be biased due to the composition of the participant pool, and further details on the recruitment process can be found in the Questionnaire Survey section.
 Modality Task Worst (%) Best (%) Average (%)
-Visual Social 46.15 66.67 54.87
-Sports 20.00 100.00 51.43
-Information Querying 53.85 71.43 62.15
-Sentiment 20.00 73.33 47.71
-Shopping 35.71 75.00 57.68
-Gaming 37.50 100.00 58.24
-Overall 35.20 81.07 55.68
-Audio Social 16.67 40.00 24.17
-Sports 20.00 42.86 33.57
-Information Querying 23.08 36.67 33.27
-Sentiment 21.43 40.00 35.36
-Shopping 23.08 60.00 44.10
-Gaming 25.00 25.00 25.00
-Overall 21.88 40.42 32.58
-Both Social 66.67 100.00 79.37
-Sports 42.86 100.00 80.95
-Information Querying 73.33 89.66 81.00
-Sentiment 86.67 93.33 89.17
-Shopping 60.00 100.00 80.00
-Gaming 57.14 87.50 71.13
-Overall 64.78 95.08 80.94
-Table D.1: Worst, Best, and Average human performance (%) across tasks and modalities.
-D.4 Quantitative GPT Evaluations
+Visual Social 66.71 94.72 79.31
+Sports 28.91 93.84 74.34
+Information Querying 77.84 95.67 89.83
+Sentiment 28.91 92.43 68.96
+Shopping 51.62 94.11 83.37
+Gaming 54.20 96.34 84.18
+Overall 28.91 96.34 80
+Audio Social 22.10 53.04 32.05
+Sports 26.52 56.83 44.51
+Information Querying 30.60 48.63 44.12
+Sentiment 28.42 53.04 46.89
+Shopping 30.60 79.56 58.48
+Gaming 33.15 33.15 33.15
+Overall 22.10 79.56 43.2
+Both (A+V) Social 78 95.81 92.86
+Sports 50.15 97.20 94.71
+Information Querying 85.80 94.91 94.77
+Sentiment 96.28 96.70 96.87
+Shopping 70.20 95.17 93.60
+Gaming 66.85 93.42 83.22
+Overall 50.15 97.20 92.67
+Table 7: Worst, Best, and Average human performance (%) across tasks and modalities.
+C.4 Quantitative GPT Evaluations
 To further assess model performance beyond multiple-choice accuracy, we conduct a quantitative evaluation of the open-ended question responses using GPT-based scoring.
-Specifically, as discussed in Sec.
-2.2, we prompt GPT-4o to act as an automated grader, evaluating the correctness and relevance of the model’s free-form answers based on predefined guidelines consistent with the task objectives.
+Specifically, as discussed in the paper, we prompt GPT-4o to act as an automated grader, evaluating the correctness and relevance of the model’s free-form answers based on predefined guidelines consistent with the task objectives in Tab.
+8 and Tab.
+9.
 This automatic assessment provides an additional perspective on the model’s generative capabilities, capturing nuances that are not reflected in standard accuracy metrics.
@@
 These findings complement the multiple-choice evaluations and expose areas where the model demonstrates strong language generation and areas where further improvement is needed.
-However, we note that these scores may be biased due to limitations in GPT-4o’s own reasoning and evaluation capabilities [28], which can introduce alignment artifacts or systematic preferences.
-Models Overall
-Human (Video) 1.03
-Human (Audio) 0.65
-Human (Video+Audio) 2.50
-w/o subs w/ subs
-Open source MLLMs w/ Truncated Videos
-InternVL2 0.896 1.055
-Qwen2.5-VL 1.651 1.929
-LLaVA-OneVision 1.774 1.124
-DeepSeek-VL2-Tiny 0.839 0.870
-DeepSeek-VL2-Small 1.128 1.471
-Closed-source MLLMs w/ Truncated Videos
-GPT-4o 1.327 2.308
-o1 2.697 3.075
-Claude 3.5 Sonnet 1.964 2.568
-Grok2-Vision 2.193 2.765
-Nova-Lite 1.444 1.859
-Nova-Pro 1.998 2.364
-Gemini 1.5 Pro 2.170 2.538
-Open source MLLMs w/ Full-Length Videos
-InternVL2 0.895 1.102
-Qwen2.5-VL 1.489 1.830
-LLaVA-OneVision 1.471 1.666
-DeepSeek-VL2-Tiny 0.675 0.808
-DeepSeek-VL2-Small 0.839 1.295
-Closed-source MLLMs w/ Full-Length Videos
-GPT-4o 1.061 2.148
-o1 2.484 2.983
-Claude 3.5 Sonnet 0.017 0.401
-Grok2-Vision 1.935 2.622
-Nova-Lite 1.653 1.708
-Nova-Pro 1.615 2.012
-Gemini 1.5 Pro 1.956 2.318
-Table D.2: GPT-4o evaluation of consistency score.
-Models Overall
-Human (Video) 1.02
-Human (Audio) 0.58
-Human (Video+Audio) 2.71
-w/o subs w/ subs
-Open source MLLMs w/ Truncated Videos
-InternVL2 0.129 1.007
-Qwen2.5-VL 1.404 1.702
-LLaVA-OneVision 1.168 1.005
-DeepSeek-VL2-Tiny 0.562 0.614
-DeepSeek-VL2-Small 1.087 1.851
-Closed-source MLLMs w/ Truncated Videos
-GPT-4o 1.277 2.103
-o1 2.484 2.803
-Claude 3.5 Sonnet 2.122 2.283
-Grok2-Vision 1.901 2.434
-Nova-Lite 1.287 1.602
-Nova-Pro 1.518 1.741
-Gemini 1.5 Pro 1.970 2.299
-Open source MLLMs w/ Full-Length Videos
-InternVL2 1.333 0.975
-Qwen2.5-VL 1.275 1.660
-LLaVA-OneVision 1.277 1.572
-DeepSeek-VL2-Tiny 0.687 1.168
-DeepSeek-VL2-Small 0.873 1.268
-Closed-source MLLMs w/ Full-Length Videos
-GPT-4o 1.277 2.103
-o1 2.484 2.803
-Claude 3.5 Sonnet 0.319 0.415
-Grok2-Vision 1.693 2.189
-Nova-Lite 1.196 1.481
-Nova-Pro 1.384 1.678
-Gemini 1.5 Pro 1.808 2.136
-Table D.3: GPT-4o evaluation of correctness score.
-Models Overall
-Human (Video) 0.90
-Human (Audio) 0.52
-Human (Video+Audio) 2.55
-w/o subs w/ subs
-Open source MLLMs w/ Truncated Videos
-InternVL2 0.586 1.120
-Qwen2.5-VL 1.746 2.033
-LLaVA-OneVision 0.953 1.090
-DeepSeek-VL2-Tiny 0.653 0.68
-DeepSeek-VL2-Small 1.162 1.493
-Closed-source MLLMs w/ Truncated Videos
-GPT-4o 1.284 2.220
-o1 3.012 3.236
-Claude 3.5 Sonnet 1.984 2.613
-Grok2-Vision 2.454 2.968
-Nova-Lite 1.339 1.701
-Nova-Pro 1.812 1.841
-Gemini 1.5 Pro 2.374 2.716
-Open source MLLMs w/ Full-Length Videos
-InternVL2 1.107 1.159
-Qwen2.5-VL 1.559 1.996
-LLaVA-OneVision 1.40 1.679
-DeepSeek-VL2-Tiny 0.839 0.953
-DeepSeek-VL2-Small 0.970 1.382
-Closed-source MLLMs w/ Full-Length Videos
-GPT-4o 1.032 1.967
-o1 2.880 3.159
-Claude 3.5 Sonnet 0.146 0.494
-Grok2-Vision 2.237 2.792
-Nova-Lite 1.273 1.562
-Nova-Pro 1.567 1.832
-Gemini 1.5 Pro 2.288 2.618
-Table D.4: GPT-4o evaluation of detailed orientation score.
-Models Overall
-Human (Video) 1.26
-Human (Audio) 0.58
-Human (Video+Audio) 2.71
-w/o subs w/ subs
-Open source MLLMs w/ Truncated Videos
-InternVL2 0.129 1.393
-Qwen2.5-VL 2.007 2.249
-LLaVA-OneVision 1.168 1.311
-DeepSeek-VL2-Tiny 0.839 0.870
-DeepSeek-VL2-Small 1.505 1.851
-Closed-source MLLMs w/ Truncated Videos
-GPT-4o 1.586 2.499
-o1 3.032 3.291
-Claude 3.5 Sonnet 2.122 2.734
-Grok2-Vision 2.468 2.960
-Nova-Lite 1.742 2.039
-Nova-Pro 2.046 2.238
-Gemini 1.5 Pro 2.435 2.801
-Open source MLLMs w/ Full-Length Videos
-InternVL2 1.332 1.366
-Qwen2.5-VL 1.818 2.187
-LLaVA-OneVision 1.725 1.995
-DeepSeek-VL2-Tiny 1.067 1.168
-DeepSeek-VL2-Small 1.292 1.655
-Closed-source MLLMs w/ Full-Length Videos
-GPT-4o 1.274 2.250
-o1 2.912 3.214
-Claude 3.5 Sonnet 0.319 0.533
-Grok2-Vision 2.240 2.751
-Nova-Lite 1.653 1.936
-Nova-Pro 1.915 2.2348
-Gemini 1.5 Pro 2.257 2.618
-Table D.5: GPT-4o evaluation of context score.
-D.5 Quantitative Evaluations by Categories
+However, we note that these scores may be biased due to limitations in GPT-4o’s own reasoning and evaluation capabilities (Li et al.
+2025a), which can introduce alignment artifacts or systematic preferences.
+Model Temporal Consistency Correctness
+Video Audio/Sub V + A/Sub Video Audio/Sub V + A/Sub
+Human 1.85 1.69 2.83 1.93 1.46 2.65
+Open-source MLLMs
+EgoGPT-7B 1.3955 1.2523 1.5634 1.2477 1.1843 1.4930
+Ola-7B 0.9178 1.2500 1.3638 0.9038 1.2171 1.3005
+VITA 1.5 0.6643 1.2160 0.7042 0.6561 1.0587 0.7007
+Qwen-2-VL 1.4554 1.3638 1.6772 1.3498 1.2254 1.6092
+Qwen-2.5-VL 1.6819 1.2793 1.9730 1.4014 1.1279 1.7477
+Qwen-2.5-Omni 1.6819 1.2793 1.9730 1.4014 1.1279 1.7477
+InternVL2 0.9742 1.3169 1.1925 0.9237 0.9143 1.1315
+LLaVA-OneVision 1.5962 1.3685 1.8052 1.4002 1.2899 1.5763
+DeepSeekVL2-small 0.9507 1.3157 1.3286 0.9953 1.1866 1.3650
+Proprietary MLLMs
+Gemini 2.0-FL 1.6526 1.5880 2.1784 1.4296 1.4390 1.9014
+Gemini 2.5-FL 1.7864 1.6244 2.1678 1.5751 1.4507 1.9167
+Claude Sonnet 3.5 2.0798 1.9155 2.6009 1.6538 1.5857 2.2300
+GPT-4o 1.5434 1.8462 2.4390 1.4167 1.5634 2.2207
+GPT-4o-mini 1.9859 1.9272 2.3850 1.5739 1.6185 2.0117
+Grok 4 2.2782 2.1854 2.8028 1.9507 1.7066 2.4354
+NOVA-Lite 0.7453 0.9167 1.0129 0.9812 1.1068 1.2066
+NOVA-Pro 0.9484 0.9930 1.3333 1.1901 1.1221 1.5387
+Table 8: GPT-4o evaluation of Temporal Consistency and Correctness scores (higher is better). Columns show performance for Video, Audio/Subtitles, and their combination.
+Model Context Detail Orientation
+Video Audio/Sub V +A/Sub Video Audio/Sub V +A/Sub
+Human 2.13 1.93 3.11 1.84 2.02 2.96
+Open-source MLLMs
+EgoGPT-7B 1.6420 1.5599 1.8732 1.2042 1.1244 1.3509
+Ola-7B 1.3146 1.6491 1.7254 1.0047 1.2805 1.3345
+VITA 1.5 0.9906 1.4918 1.0610 0.7465 1.2735 0.7958
+Qwen-2-VL 1.7570 1.6549 2.0505 1.4343 1.4178 1.7430
+Qwen-2.5-VL 1.9824 1.6197 2.2829 1.7336 1.3885 2.0669
+InternVL2 1.3462 1.3521 1.5164 1.1303 1.0505 1.2523
+LLaVA-OneVision 1.8662 1.7136 2.0129 1.5129 1.4284 1.7031
+DeepSeekVL2-small 1.3920 1.6984 1.7418 1.0340 1.5998 1.3580
+Proprietary MLLMs
+Gemini 2.0-FL 1.8204 1.8439 2.1784 1.5880 1.5293 2.0117
+Gemini 2.5-FL 1.9930 1.9120 2.3216 1.7653 1.6678 2.0692
+Claude Sonnet 3.5 2.1725 2.0493 2.6960 2.0129 1.9096 2.5575
+GPT-4o 1.7570 2.0634 2.6103 1.4178 1.7946 2.3169
+GPT-4o-mini 2.1444 2.1397 2.5716 1.7277 1.8662 2.2336
+Grok 4 2.5141 2.2488 2.9636 2.4660 2.1185 2.9566
+NOVA-Lite 1.4178 1.5516 1.6843 1.1749 1.3451 1.3955
+NOVA-Pro 1.7312 1.5387 2.0610 1.5282 1.3439 1.8204
+Table 9: GPT-4o evaluation of Context score and Detail Orientation score (higher is better). Columns show performance for Video only, Audio/Subtitles only, and Video+Audio/Subtitles.
+C.5 Quantitative Evaluations by Categories
 To provide a comprehensive understanding of model performance across the diverse challenge dimensions of the benchmark, we report quantitative results for all probed models across each defined split.
-This includes evaluations broken down by agentic task categories, difficulty levels, audio characteristics, topics, and input modalities.
+This includes evaluations broken down by agentic task categories, difficulty levels, audio characteristics, topics, and input modalities, as shown in Tab.
+10.
 These category-wise results allow us to examine how different models handle specific functional scenarios, content types, and levels of complexity.
@@
 Together, these fine-grained evaluations provide a detailed view of model behavior across the full spectrum of tasks and shed light on future work on multimodal understanding and generalization.
-GPT-4o o1 Gemini 1.5 Pro Nova Pro Grok2 Claude 3.5 Sonnet
-Difficulty Easy 69.8 79.3 69.5 51.5 65.1 68.9
-Medium 67.0 76.0 69.4 47.5 61.2 64.6
-Hard 49.6 67.4 49.6 34.1 47.4 44.4
-Agentic Tasks Social 64.6 69.7 62.6 49.5 60.6 66.7
-Sentiment 56.8 74.1 66.7 46.9 51.9 51.9
-Egocentric Agent 62.7 69.5 57.6 40.7 54.2 64.4
-Information Querying 67.5 75.5 69.4 47.2 62.6 64.9
-Sports 66.4 79.0 64.7 40.3 58.8 56.3
-Gaming 64.1 72.8 58.3 54.4 52.4 59.2
-Shopping 68.3 85.7 74.6 47.6 73.0 73.0
-Audio Type Natural Sound 57.4 66.2 54.4 35.3 50.0 48.5
-Speech 67.7 76.7 69.3 44.8 61.7 65.0
-Music 57.7 61.5 69.2 38.5 57.7 57.7
-Artificial Sound 70.0 78.0 56.0 58.0 54.0 60.0
-Mixed Sounds 63.4 78.2 66.0 52.3 63.0 64.9
-Topics Humanities and Society 71.4 78.9 77.4 51.1 66.9 72.2
-Geography and Travel 56.2 68.8 50.0 37.5 43.8 43.8
-Technology and Gaming 65.6 71.8 61.1 50.4 53.4 57.3
-Science and Knowledge 60.7 64.3 53.6 28.6 50.0 53.6
-Movies, TV and Animations 55.2 68.7 55.2 40.3 43.3 55.2
-Arts and Performance 51.7 65.5 58.6 41.4 58.6 75.9
-Pets and Animals 62.2 68.9 64.4 48.9 62.2 57.8
-Business and Commerce 72.1 86.0 79.1 60.5 76.7 75.6
-Life and Practical Skills 74.0 79.8 77.9 51.9 71.2 69.2
-Sports and Adventure 63.5 80.2 65.1 46.0 57.1 57.9
-Social Trends and Reactions 55.6 77.8 57.4 33.3 61.1 59.3
-Vehicles and Transportation 88.9 100.0 55.6 55.6 55.6 55.6
-Low Quality and Extended Content 54.2 58.3 37.5 16.7 50.0 54.2
-Multimodal Abilities Visual Reasoning 59.2 73.2 64.5 42.1 57.5 59.6
-Emotional Inference 60.5 69.8 62.8 46.5 53.5 48.8
-Situational Reasoning 65.0 80.0 70.0 40.0 75.0 75.0
-Causal Reasoning 65.7 73.1 67.2 50.7 58.2 68.7
-Factual Recall 68.9 79.2 68.2 49.2 63.2 65.0
-Spatial Reasoning 62.1 62.1 58.6 41.4 44.8 55.2
-Relationship 75.0 81.2 62.5 50.0 81.2 75.0
-Temporal Distance 66.7 66.7 58.3 58.3 50.0 66.7
-Overall 65.4 75.9 66.3 46.9 60.6 63.1
-Table D.6: Performance comparison across different models and categories on truncated videos.
-GPT-4o o1 Gemini 1.5 Pro Nova Pro Grok2 Claude 3.5 Sonnet
-Difficulty Easy 64.8 73.0 69.4 58.2 59.2 17.4
-Medium 62.2 72.4 68.0 58.1 55.2 14.8
-Hard 46.5 65.4 55.1 41.7 53.5 17.3
-Agentic Tasks Social 66.7 65.6 59.4 63.5 62.5 9.4
-Sentiment 54.3 67.9 64.2 46.9 55.6 24.7
-Egocentric Agent 55.3 61.7 55.3 19.1 34.0 14.9
-Information Querying 63.9 75.4 68.3 60.7 57.9 17.5
-Sports 55.3 69.9 66.0 55.3 59.2 14.6
-Gaming 60.0 71.8 64.7 48.2 48.2 22.4
-Shopping 60.4 75.7 76.6 64.0 62.2 10.8
-Audio Type Natural Sound 70.5 79.8 61.5 53.8 43.1 30.8
-Speech 56.2 81.2 70.3 57.3 60.7 14.4
-Music 55.3 70.2 56.0 48.0 56.0 16.0
-Artificial Sound 51.9 59.3 58.0 48.0 48.0 28.0
-Mixed Sounds 58.2 59.7 63.8 55.0 54.6 12.8
-Topics Humanities and Society 57.1 67.9 79.1 66.7 64.3 27.1
-Geography and Travel 60.0 68.9 68.8 56.2 56.2 18.8
-Technology and Gaming 61.4 72.9 62.3 50.9 50.0 17.5
-Science and Knowledge 69.4 75.3 63.0 55.6 51.9 11.1
-Movies, TV and Animations 55.0 68.8 59.7 52.2 46.3 11.9
-Arts and Performance 57.7 75.0 57.1 46.4 57.1 17.9
-Pets and Animals 44.4 88.9 60.0 57.8 57.8 17.8
-Business and Commerce 62.5 62.5 77.1 65.7 67.1 12.9
-Life and Practical Skills 57.8 68.6 64.7 57.6 62.4 8.2
-Sports and Adventure 43.9 65.9 69.7 52.3 55.0 16.5
-Social Trends and Reactions 72.2 77.8 57.7 51.9 51.9 15.4
-Vehicles and Transportation 53.4 63.8 44.4 44.4 33.3 0.0
-Low Quality and Extended Content 63.3 73.9 50.0 20.8 50.0 8.3
-Multimodal Abilities Visual Reasoning 62.5 66.7 64.2 51.5 54.4 14.7
-Emotional Inference 68.8 75.0 58.5 53.7 43.9 19.5
-Situational Reasoning 81.8 90.9 66.7 55.6 77.8 16.7
-Causal Reasoning 53.8 56.9 65.5 55.2 55.2 25.9
-Factual Recall 63.8 71.9 69.0 58.1 58.8 15.9
-Spatial Reasoning 52.0 60.0 45.8 41.7 41.7 16.7
-Relationship 66.0 82.0 75.0 68.8 68.8 0.0
-Temporal Distance 56.4 73.9 81.8 54.5 45.5 18.2
-Overall 60.6 71.5 66.5 55.5 56.5 16.3
-Table D.7: Performance comparison across different models and categories on full-length videos.
-D.6 Common Error Modes
-To better understand the limitations of the evaluated models, we analyze common error patterns observed across the benchmark.
+GPT-4o Gemini 2.5 FL Qwen 2.5 Omni Nova Lite Nova Pro Ola-7B
+Difficulty
+Easy 66.7 59.8 52.1 55.0 62.5 57.4
+Medium 67.4 54.9 48.8 52.4 54.9 54.1
+Hard 49.7 43.0 45.2 38.4 43.0 41.1
+Agentic Tasks
+Social 62.6 55.6 47.5 48.5 50.5 52.5
+Sentiment 56.8 54.3 58.0 42.0 55.6 55.6
+Egocentric Agent 59.3 39.0 42.4 44.1 54.2 57.6
+Information Querying 68.3 60.0 45.3 52.1 54.7 49.1
+Sports 60.5 47.1 49.6 50.4 58.0 52.9
+Gaming 61.2 49.5 41.7 39.8 47.6 46.6
+Shopping 68.3 61.9 64.3 69.0 67.5 63.5
+Audio Type
+Natural Sound 52.9 51.5 52.9 35.3 45.6 48.5
+Speech 66.4 56.7 47.8 53.8 57.6 52.0
+Music 61.5 50.0 53.8 42.3 57.7 38.5
+Artificial Sound 64.0 52.0 48.0 48.0 50.0 58.0
+Mixed Sounds 63.0 53.1 51.5 51.5 56.1 56.5
+Topics
+Humanities and Society 71.4 64.7 56.4 54.1 63.9 54.1
+Geography and Travel 56.2 50.0 37.5 56.2 31.2 56.2
+Technology and Gaming 65.6 55.0 42.0 43.5 48.9 48.9
+Science and Knowledge 60.7 50.0 50.0 46.4 35.7 60.7
+Movies, TV and Animations 53.7 37.3 34.3 43.3 52.2 38.8
+Arts and Performance 51.7 44.8 51.7 44.8 44.8 48.3
+Pets and Animals 57.8 66.7 55.6 40.0 53.3 53.3
+Business and Commerce 69.8 66.3 68.6 73.3 69.8 66.3
+Life and Practical Skills 72.1 60.6 51.9 60.6 61.5 61.5
+Sports and Adventure 61.1 49.2 48.4 46.0 58.7 50.0
+Social Trends and Reactions 53.7 48.1 44.4 46.3 40.7 50.0
+Vehicles and Transportation 66.7 33.3 55.6 44.4 77.8 66.7
+Low Quality and Extended Content 58.3 29.2 25.0 41.7 50.0 37.5
+Multimodal Abilities
+Visual Reasoning 58.3 54.4 45.6 50.0 51.8 52.6
+Emotional Inference 53.5 65.1 51.2 39.5 44.2 44.2
+Situational Reasoning 70.0 65.0 35.0 55.0 65.0 70.0
+Causal Reasoning 67.2 49.3 44.8 49.3 55.2 47.8
+Factual Recall 67.3 55.4 52.9 53.8 58.6 54.2
+Spatial Reasoning 55.2 41.4 48.3 34.5 55.2 37.9
+Relationship 68.8 50.0 50.0 56.2 62.5 75.0
+Temporal Distance 75.0 50.0 50.0 41.7 50.0 58.3
+Overall 64.0 54.7 49.5 50.9 55.8 53.1
+Table 10: Performance comparison across different models and categories with Video + Audio/Subtitle modalities.
+C.6 Common Error Modes
+To better understand the limitations of the state of the art models, we analyze common error patterns observed across the benchmark.
 We find that model failures often arise from modality-specific weaknesses, such as misinterpretation of complex visual cues, inability to process nuanced audio signals, or challenges in aligning multimodal information with nuanced context.
+For illustration, we highlight representative failure cases from OpenAI o1 and Gemini 1.5 Pro, chosen because they are among the strongest available models; however, the same error patterns recur across the models we evaluated.
 Common failure modes are showcased in Fig.
-D.2,  D.3,  D.4,
-Figure D.2: Error analysis showing that o1 fails to correctly answer the question when the audio cannot be transcribed into text-based subtitles, leading to an incorrect connection between the cityscape video and the instrumental music.
-Figure D.3: Error analysis showing inconsistent responses from o1, where it failed to accurately identify Hika-chan, when the question was rephrased slightly.
-Figure D.4: Error analysis highlighting o1’s failure in correctly synchronizing the audio with visual cues, leading to misinterpretation of the woman picking up tree ornaments.
-Figure D.5: Error analysis highlighting o1’s contextual misinterpretation, attributing the man’s smile to a general amusing memory rather than the dog’s barking resembling gunfire.
+5,  6,  7,  8,  9,  10.
+Figure 5: Error analysis showing that o1 fails to correctly answer the question when the audio cannot be transcribed into text-based subtitles, leading to an incorrect connection between the cityscape video and the instrumental music.
+Figure 6: Error analysis showing inconsistent responses from o1, where it failed to accurately identify Hika-chan, when the question was rephrased slightly.
+Figure 7: Error analysis highlighting o1’s failure in correctly synchronizing the audio with visual cues, leading to misinterpretation of the woman picking up tree ornaments.
+Figure 8: Error analysis highlighting o1’s contextual misinterpretation, attributing the man’s smile to a general amusing memory rather than the dog’s barking resembling gunfire.
 Additionally, Gemini Pro 1.5 exhibits visually induced audio hallucination, mistaking the dog’s open-mouth appearance for barking and associating it with a crying baby.
-Figure D.6: Error analysis showing o1 and Gemini Pro 1.5 both struggling with spatial misjudgment and temporal misalignment, leading to errors in estimating the number of defenders due to difficulties in distinguishing actual close defenders and potential misalignment with the described action.
-Figure D.7: Error analysis showing Gemini Pro 1.5 providing inconsistent responses despite clear visuals and audio cues, failing to accurately identify the cause of the glass breaking due to rephrased question structure.
-Appendix E Questionnaire Survey
+Figure 9: Error analysis showing o1 and Gemini Pro 1.5 both struggling with spatial misjudgment and temporal misalignment, leading to errors in estimating the number of defenders due to difficulties in distinguishing actual close defenders and potential misalignment with the described action.
+Figure 10: Error analysis showing Gemini Pro 1.5 providing inconsistent responses despite clear visuals and audio cues, failing to accurately identify the cause of the glass breaking due to rephrased question structure.
+Appendix D D. Questionnaire Survey
 To establish human performance baselines for comparison with the probed models, we conducted a questionnaire study approved by our institutional IRB.
@@
 Participants answered questions presented in the same format as those used for model evaluation.
-The survey included a total of 200 questions curated from the dataset, covering diverse subcategories within each defined split.
-For each participant, the questionnaire consisted of two parts: one multiple-choice question (MCQ) with eight answer options, followed by one open-ended question from a different video.
+The survey included a total of 261 questions curated from the dataset, covering diverse subcategories within each defined split.
+For each participant, the questionnaire consisted of either one multiple-choice question (MCQ) with eight answer options, or an open-ended question from a different video.
 To minimize content overlap and reduce potential bias, no video was reused between questions within a single questionnaire.
 This approach ensured broad coverage of the dataset’s taxonomy while avoiding contamination across tasks.
+We recruited 382 participants through Amazon Mechanical Turk service with approval rate >>96% to answer the questions with A, S, V, A+V, S+V as inputs with media from MAVERIX and MAVERIX-Long.
+Both MCQ responses and open-ended responses are recorded.
 For evaluation, MCQ responses were scored against the ground truth to calculate accuracy, while open-ended answers were assessed using the same GPT-4o evaluation pipeline applied to model outputs, ensuring consistency across human and model scoring.
@@
 The survey interfaces for the conditions are shown in Fig.
-E.2 for audio only, Fig.
-E.3 for visual only, Fig.
-E.1 for both modalities.
+12 for audio only, Fig.
+13 for visual only, Fig.
+11 for both modalities.
 To prevent cross-condition contamination and ensure focused evaluation, each participant completed the entire study under only one selected modality.
 This human study provides us with valuable reference points for interpreting model results and understanding how humans perform under comparable multimodal constraints.
-Figure E.1: Screenshot of the questionnaire under the audio-plus-visual condition, showcasing the interface used to assess human performance across multiple-choice questions.
+Figure 11: Screenshot of the questionnaire under the audio-plus-visual condition, showcasing the interface used to assess human performance across multiple-choice questions.
 Results offer a baseline for comparison with multimodal model performance across diverse tasks and difficulty levels.
-Figure E.2: Questionnaire screenshot showing open-ended questions under the audio-only condition.
+Figure 12: Questionnaire screenshot showing open-ended questions under the audio-only condition.
 Responses were evaluated using the GPT-4o grading pipeline, capturing the quality of language generation based solely on audio context.
-Figure E.3: Questionnaire screenshot showing open-ended questions under the visual-only condition.
+Figure 13: Questionnaire screenshot showing open-ended questions under the visual-only condition.
 GPT-4o evaluation scores assess human understanding based solely on visual information, without the aid of audio context.
-Appendix F Dataset Distribution Continued
+Appendix E E. Dataset Distribution Continued
 In this section, we provide a detailed breakdown of the taxonomy distributions within each defined agentic task category.
@@
 The corresponding distributions are visualized in Fig.
-F.1 for Gaming, Fig.
-F.2 for Egocentric Agent, Fig.
-F.3 for Information Querying, Fig.
-F.4 for Sentiments, Fig.
-F.5 for Shopping, Fig.
-F.6 for Social, and Fig.
-F.7 for Sports.
+14 for Gaming, Fig.
+15 for Egocentric Agent, Fig.
+16 for Information Querying, Fig.
+17 for Sentiments, Fig.
+18 for Shopping, Fig.
+19 for Social, and Fig.
+20 for Sports.
 These visualizations offer insight into the internal structure of each agentic task and highlight the variety of situations represented within the dataset.
-Figure F.1: Taxonomy composition in Gaming.
-Figure F.2: Taxonomy composition in Egocentric Agent.
-Figure F.3: Taxonomy composition in Information Querying.
-Figure F.4: Taxonomy composition in Sentiments.
-Figure F.5: Taxonomy composition in Shopping.
-Figure F.6: Taxonomy composition in Social.
-Figure F.7: Taxonomy composition in Sports.
+Figure 14: Taxonomy composition in Gaming.
+Figure 15: Taxonomy composition in Egocentric Agent.
+Figure 16: Taxonomy composition in Information Querying.
+Figure 17: Taxonomy composition in Sentiments.
+Figure 18: Taxonomy composition in Shopping.
+Figure 19: Taxonomy composition in Social.
+Figure 20: Taxonomy composition in Sports.